### v1

# Revisiting MAE pre-training for 3D medical image segmentation

Abstract
Self-Supervised Learning (SSL) presents an exciting opportunity to unlock the potential of vast, untapped clinical datasets, for various downstream applications that suffer from the scarcity of labeled data. While SSL has revolutionized fields like natural language processing and computer vision, their adoption in 3D medical image computing has been limited by three key pitfalls: Small pre-training dataset sizes, architectures inadequate for 3D medical image analysis, and insufficient evaluation practices. We address these issues by i) leveraging a large-scale dataset of 44k 3D brain MRI volumes and ii) using a Residual Encoder U-Net architecture within the state-of-the-art nnU-Net framework. iii) A robust development framework, incorporating 5 development and 8 testing brain MRI segmentation datasets, allowed performance-driven design decisions to optimize the simple concept of Masked Auto Encoders (MAEs) for 3D CNNs. The resulting model not only surpasses previous SSL methods but also outperforms the strong nnU-Net baseline by an average of approximately 3 Dice points. Furthermore, our model demonstrates exceptional stability, achieving the highest average rank of 2 out of 7 methods, compared to the second-best method’s mean rank of 3. Our code is made available here.
1 Introduction
In recent years, the concept of Self-Supervised Learning (SSL) has emerged as a driving factor in data-rich domains, enabling large-scale pre-training that facilitates the learning of robust and transferable general-purpose representations (Assran et al., 2023; Oquab et al., 2023; He et al., 2022). This paradigm shift has been instrumental in advancing various fields, particularly in domains with abundant labeled data like NLP or natural vision. In the domain of 3D medical image computing, this trend has not caught on.
Currently, the domain is either focused on training from-scratch, mainly using the nnU-Net framework by Isensee et al. (2021), or using supervised pretraining, which is limited by the cost associated with annotated data (Wasserthal et al., 2023; Ulrich et al., 2023; Huang et al., 2023). The usage of supervised pretraining implies the willingness of the domain to adopt pretraining and calls to question the currently established SSL methods in the domain. We believe this lack of widespread adoption of previously established methods can be attributed to three major pitfalls of previous SSL research in this field:
P1 - Limited dataset size: Many SSL approaches have been developed on very few unlabeled volumes, often being trained on fewer than 10,000 images Zhou et al. (2021); Wu et al. (2024); Tang et al. (2024; 2022); Zhuang et al. (2023); Hatamizadeh et al. (2021), almost approaching scales of supervised dataset sizes. These datasets tend to be pooled from publicly available annotated datasets, as larger datasets pose a greater hurdle to acquire. While many hospitals possess 3D medical images in the millions, they are locked away from the public due to patient privacy concerns. While some larger open-source datasets exist, e.g., the Adolescent Brain Cognitive Decline (ABCD) dataset of the NIH (N=40k) or the UK Biobank (UK-BB) (N=120k), they restrict access pending an internal review board’s approval, posing a hurdle for open science. More recently, the UK-BB stopped allowing downloading data to local hardware, signaling a public decrease in the community’s willingness to share data.
P2 - Outdated Backbones: Many studies develop their SSL method on non-state-of-the-art architectures, e.g., utilizing transformers (Wu et al., 2024; Tang et al., 2022; Wang et al., 2023; Chen et al., 2023). While transformers are prevalent in the 2D natural imaging domain (Dosovitskiy, 2020), recent architectures leveraging attention (Vaswani, 2017) have so far not been able to reach state-of-the-art performance in 3D medical segmentation. In fact, well-configured 3D U-Net (Çiçek et al., 2016; Ronneberger et al., 2015) inspired CNNs dominate 3D medical image segmentation, outperforming transformer-based models by a large margin Isensee et al. (2024). This underscores the need for SSL methods that can be seamlessly integrated with CNNs to harness their full potential in medical image analysis on downstream tasks.
P3 - Insufficient Evaluation: Existing methods often lack rigorous evaluation, masking the methods’ efficacy (or lack thereof). This is represented through: i) Evaluating on too few datasets to show generalization of the pre-training ii) Stacking multiple contributions, e.g., novel architecture with new pre-training, which does not allow one to draw conclusions if the pre-training is effective without the architecture (Wang et al., 2023) iii) Comparing to bad from-scratch baselines, like a badly configured, outdated Çiçek et al. (2016) model instead of a well-configured 3D nnU-Net CNN baseline (Isensee et al., 2021). iv) Evaluating their method on seen data they pre-trained on. We want to emphasize that we do not intend to point fingers, but to raise awareness that evaluation matters and insufficient evaluation can lead to a lack of clarity which methods are the best. This observation is similar to the recent study by Isensee et al. (2024), that shows that this is also prevalent in the field of medical image segmentation when training from-scratch.
In this paper, we carefully avoided all these pitfalls while exploring the Masked Auto Encoder (MAE) paradigm for 3D CNNs, with the recent adaptations introduced by Tian et al. (2023); Woo et al. (2023), and highlight that MAEs are exceeding the current state-of-the-art SSL methods in 3D medical image segmentation given proper configuration. Our contributions can be summarized as follows:
  1. 1.
    We evade Pitfall 1 by leveraging a dataset collection of 44k 3D MRI volumes to develop our self-supervised pretraining, exceeding the majority of 3D medical image segmentation SSL methods in scale (Zhou et al., 2021; Wu et al., 2024; Tang et al., 2024; 2022; Zhuang et al., 2023; Hatamizadeh et al., 2021).
  2. 2.
    We evade Pitfall 2 by utilizing the state-of-the-art Residual Encoder U-Net CNN architecture from Isensee et al. (2024) as the backbone. Moreover, we use this backbone for all baseline SSL methods - allowing us to quantify the utility of the pre-training scheme for pre-training CNNs.
  3. 3.
    We evade Pitfall 3 by employing 5 development and 7 testing datasets, spanning a diverse set of downstream targets. This includes head and neck organs and pathologies, datasets with novel image modalities not seen during pre-training, and datasets of the same pathology acquired at different centers.
By subsequently optimizing MAEs for CNNs based on performance on the development pool, we propose the novel sparse MAE inspired pre-training paradigm, unleashing the potential of huge amounts of not annotated data in the medical domain. We show that the choice of the best fine-tuning strategy is crucial. Moreover, we evaluate our method across a range of critical 3D medical scenarios, including low-data regimes, accelerated fine-tuning schedules, and generalization across unseen centers and modalities.
Figure 1: Overview of the model development pipeline. During pretraining, we carefully develop our method with 44k images (P1), with a state-of-the-art residual encoder U-Net (P2) and evaluate it on dive development datasets (P3). In the testing phase we utilize a total of thirteen datasets (P3) to verify the utility of our developed method.
2 Development Framework
The goal of this paper is to develop a robust SSL pretraining method. Due to the limited prior work in the 3D medical domain, many design choices need to be made. We address this by sequentially validating each methodological contribution on five downstream development datasets before testing the final configuration on eight untouched test datasets. To reduce the search space and disentangle the effects of SSL pre-training from other basic design choices, we choose to keep some parameters fixed based on best practices in the domain:
(i) The architecture used is always the same state-of-the-art residual encoder U-Net architecture (ResEnc U-Net) (Isensee et al., 2024). (ii) The input patch size is [160x160x160]. (iii) All images are resampled to the target spacing of [1x1x1] mm3 (Roy et al., 2023). (iv) All images are z-score normalized to zero mean and unit variance (Isensee et al., 2021). (v) As optimizer, we use SGD with a decreasing poly-learning rate (Chen et al., 2018) following nnU-Net (Isensee et al., 2021). (vi) We always employ random sampling during pre-training, irrespective of the prevalence of the different MR modalities.
Pre-training Dataset
To develop our pre-training method, we utilize a proprietary brain MRI dataset sourced from over 44 centers containing over 9k patients comprising a total of approx. 44k 3D MRI scans. Due to the variety of data sources, this dataset contains images from more than 10 different MR scanners, various MR modalities and a diverse patient population. For more details on the distributions we refer to Figure 2.
Since this data is sourced directly from clinical examinations, it includes empty or broken images, poor-quality images and so-called scout scans used to determine the field of view of the patient in the MR. Since these scans are not used in diagnostics, these images are filtered by discarding (a) images with a field of view of $<50mm$ in any axis, (b) images with a spacing $>6.5mm$ in any direction and (c) images of file size $<200kb$, which indicate an empty image. Moreover due to the low quantity of MR Angiography, Susceptibility weighted Images (SWI) and Proton Density (PD) weighted images, we restrict our training data to only include T1, T2, T1 FLAIR and T2 FLAIR images, resulting in our final pre-training dataset of $39,168$ MR images.
Development Datasets
After pre-training we fine-tune on five datasets and calculate the average Dice Similarity Coefficient (DSC) per dataset to evaluate the effectiveness of the pre-training. The multiple datasets are essential to ensure that our design choices do not overfit to a specific MRI modality or pathological target. Specifically, we utilize:
  1. 1.
    MS FLAIR (D1): Consensus delineations of multiple sclerosis (MS) lesions on T2-weighted FLAIR images(Muslim et al., 2022).
  2. 2.
    Brain Mets (D2): Brain metastases imaged through T1, contrast enhanced gradient echo T1ce, contrast enhanced spin echo T1 and a T2 FLAIR sequence acquired in the Stanford University Hospital (Grøvik et al., 2020).
  3. 3.
    Hippocampus (D3): The Hippocampus dataset, task 4 of the medical segmentation decathlon (MSD) (Antonelli et al., 2022), contains delineations of the anterior and posterior Hippocampus in T1 weighted MRI (Simpson et al., 2019).
  4. 4.
    Atlas22 (D4): Anatomical Tracings of Lesions After Stroke (ATLAS) on T1 weighted images. We use the Atlas R2.0 dataset from Liew et al. (2022).
  5. 5.
    CrossModa (D5): Delineations of intra-meatal and extra-meatal vestibular schwannoma tumors and cochlea delineations in contrast enhanced T1 weighted MRI (Dorent et al., 2023).
Of all these datasets, we set aside a hold-out test set comprising 20% of all images before start of the method development process. The remaining images were further split 80/20 into training and validation sets for the development process.
Test Datasets
Additionally, eight hold-out test sets were used to evaluate the efficacy of our learned representations when fine-tuning them to segment other target structures.
  1. 1.
    Cosmos (D6): This dataset features carotid vessel wall segmentation and atherosclerosis diagnosis, of which we use the contours to evaluate segmentation performance (Chen et al., 2022).
  2. 2.
    HaNSeg (D7): This dataset contains segmentations of 30 organs at risk (OAR), with associated T1 MRI and CT, of which we only use the MR images for model development (Podobnik et al., 2023).
  3. 3.
    Isles22 (D8): This dataset contains annotations of ischemic stroke lesions, with associated diffusion-weighted imaging (DWI), apparent diffusion coefficient (ADC) and T2 FLAIR (Hernandez Petzsche et al., 2022).
  4. 4.
    HNTS-MRG24(D9): T2-weighted MR images of pre-treatment oropharyngeal cancer and metastatic lymph nodes with associated annotations (Wahid et al., 2024).
  5. 5.
    BraTS Africa (D10): This dataset contains 1.5 Tesla T1, T1ce, T2, and T2 FLAIR MRIs of glioblastoma and high-grade gliomas imaged in Nigeria.
  6. 6.
    T2 Aneurysms (D11): This proprietary dataset contains 240 T2 MRI images of segmented brain aneurysms with the surrounding brain tissue.
  7. 7.
    TOF Angiography Aneurysms (D12): This proprietary dataset contains 144 time-of-flight MR-angiography images of segmented brain aneurysms with the surrounding brain tissue.
  8. 8.
    BraTS Mets (D13): This dataset holds brain metastases segmentations on T1ce MRIs, similar to D2. However, instead of fine-tuning on it we use it to measure generalization when inferring models trained on D2 on it.
For all test datasets that are fine-tuned on (D6-D11) we use an 80/20 split for fine-tuning and for testing, as we fine-tune each method only once without any interventions. For D13 all data is used, as no training is conducted.
[t]
Figure 2: Hierarchical composition of the nnSSL foundation model’s pre-training dataset. (a) Geographic distribution of 44 participating centers. (b) Patient demographics (age, sex, weight). (c) Scanner specifications (magnetic field strength, manufacturer, scanner) of examinations. (d) Acquisition parameters (orientation, contrast administration, weighting) of image volumes.
3 Revisiting 3D MAEs
Configuration D1 D2 D3 D4 D5 Avg. D1-D5
Base 49.96 72.62 89.03 63.45 81.67 71.35
+ Sparse Conv. + BN. 50.34 73.34 88.91 62.96 81.25 71.36
+ w/ Mask Token 50.02 73.21 88.92 62.84 81.83 71.37
+ w/ Dens. Conv. 51.02 74.07 88.91 62.81 81.50 71.66
(a) Sparsification: Introducing all sparsification adaptations showed best results, with most gains from adding the additional densification conv layer between encoder and decoder.
Mask ratio D1 D2 D3 D4 D5 Avg. D1-D5
30 % 50.25 71.37 88.90 63.18 81.70 71.08
45 % 50.60 70.83 88.97 63.27 81.73 71.08
60 % 50.62 73.56 88.96 63.37 81.48 71.60
75 % 51.02 74.07 88.91 62.81 81.50 71.66
90 % 50.56 72.51 88.98 62.41 81.49 71.19
U[60%-90%] 51.49 74.01 88.83 62.39 81.54 71.65
(b) Masking ratio: Masking ratios between 60% and 75% worked best when choosing static ratios with a dynamic range including higher masking ratios performing the best.
Pretraining D1 D2 D3 D4 D5 Avg. D1-D5
No Dyn. 45.56 72.26 88.80 60.44 82.61 69.93
No Fixed 49.37 69.13 88.78 60.74 81.33 69.87
VoCo 50.35 67.20 88.22 57.82 80.29 68.77
VF 49.93 69.58 88.83 61.75 81.48 70.31
MG 50.50 71.14 88.83 63.29 82.15 71.18
S3D-B (ours) 51.49 74.01 88.83 62.39 81.54 71.65
S3D-L (ours) 51.42 72.84 89.09 63.30 82.15 71.76
(c) Compound scaling: When comparing our scaled S3D-L, trained for more iterations with higher batch size and lr, only a modest increase in performance can be observed over the much more compute-efficient S3D-B. Moreover, when comparing models on the same scale S3D-B exceeds all baselines on our development validation datasets.
Table 1: Development Experiments: During development, we evaluate the Average DSC on all development datasets to quantify the best configuration. Methods of the same configuration are denoted through common colors.
Masked autoencoders (MAEs) are a well-established pre-training paradigm in the natural imaging domain and in the medical image segmentation domain for transformers. In this section, we investigate this paradigm and optimize it for 3D medical image segmentation using a ResEnc U-Net architecture of Isensee et al. (2024).
Default parameters
MAEs are trained by masking an input image to a certain degree and training the network to reconstruct the occluded regions, minimizing deviations betweem reconstruction and original image. In our experiments, we train the MAE with a L2-Loss in the z-score normalized voxel space and only calculate the reconstruction loss where regions were masked. Moreover, we do not remove skip-connections, following the general consensus of Woo et al. (2023), Tian et al. (2023), and He et al. (2022). The default hyperparameters (as used by the model denoted in gray in Table 1) are learning rate 1e-2, weight decay 3e-5, batch size 6, SGD optimizer with Nesterov momentum 0.99, masking ratio of 75% trained with a PolyLR schedule for 250k steps (this represent 1000 epochs in the nnU-Net framework) and minor spatial augmentations of affine scaling, rotation and mirroring.
Sparsification
When masking the input image, CNNs are not able to ignore the masked regions in the same manner as transformers can. To address this, Tian et al. (2023) proposed to adapt the CNN architectures to better fit the sparse inputs: (a) Sparse Convolutions and Normalization:Through the receptive field of convolutions masked-out regions are iteratively eroded from their boundaries. By re-applying the masked regions after every convolution this problem can be resolved. Moreover, the masks can introduce a problematic shift in the normalization layer statistics due to the introduced zero values. To resolve this normalization is constrained to only consider the non-masked values. (b) Mask Token:Instead of feeding the light-weight decoder feature maps with zeroed mask regions, the regions are densified by filling them with a learnable mask token, simplifying the reconstruction task of the decoder. (c) Densification Convolution:After filling the masks with the Mask Token and before passing the feature maps to the decoder, a [3x3x3] convolution is applied to the feature maps at every resolution except the highest resolution to prepare the representations for decoding.
Results of these changes are visualized in Table 1(a). The adaptations are introduced iteratively, meaning the ’MaskToken’ ablation is only applied together with the Sparse Convolutions and Normalization. It can be observed that the full set of adaptations improves performance by an average of $0.3$ DSC points across our development datasets. Subsequently, all adaptations are kept and the following evaluations are presented with these changes applied.
Masking strategy
The masked region is determined by sampling randomly in the CNN’s bottleneck of shape [5x5x5] and up-sampling these regions to the input resolution, to ensure the masks align in the bottleneck of the CNN architectures. This results in masking regions of [32x32x32] voxels of non-overlapping regions in the input. As a sampling strategy, we follow random masking, as previous work showed no benefit of structured masking for images nor videos (He et al., 2022; Feichtenhofer et al., 2022). In the scope of the development phase we explore 5 static masking ratios between 30 and 90%, and evaluate a dynamic masking ratio randomly masking between 60% and 90%.
Results are presented in Table Table 1(b) and highlight that the masking ratios of 60%, 75%, and the dynamic masking ratio of 60% to 90% perform equally well. Due to the highly similar performance, we choose to proceed with a dynamic masking ratio over the static masking ratio, due to expecting this masking to be more difficult to learn and with the upcoming scaling experiment in mind. We refer to this model as Spark3D-Base (S3D-B= going forward.
Scaling
MAEs are known to benefit from scaling. We evaluate the effect of compound scaling by increasing the batch size by x8 to $48$, the learning rate to $23e-$ 2, and the iterations by x4 to $1M$. We refer to this model as S3D-Large (S3D-L) to denote the higher compute resources, but note that the architecture remains identical, to maintain easy adaptation of the parameters. Results are presented in Table 1(c). It can be observed that this x32 increase in compute only resulted in a slight increase in performance of 0.1 DSC points.
Fine-tuning strategy
Given a pre-trained model, a crucial question arises: Which weights to transfer and how to schedule the fine-tuning? We investigated various different schedules. Regarding weight transfer we investigated transferring (i) both the encoder and decoder [Uncaptioned image], or (ii) only the encoder [Uncaptioned image] with a randomly initialized decoder. Regarding the fine-tuning schedule, we investigate whether to use a learning rate warm-up of 12.5k steps, ramping the learning rate up to the maximum LR. When only transferring the encoder, an additional warm-up of only the decoder is investigated to adapt the randomly initialized decoder to the pre-trained encoder. In some configurations, this results in two learning rate warm-ups of 12.5k steps each. Additionally, we investigate whether to keep the encoder frozen [Uncaptioned image] for the entire fine-tuning process, or to fine-tuning encoder weights as well. Lastly, we investigate whether to decrease the peak learning rate to $31e-$ 3, $41e-$ 4 or to keep it at the default of $21e-$ 2.
Results are presented in Table 2 and allow three important observations to be made: (i) Warm-up stages are essential:Not applying a warm-up step significantly reduces performance. Including a warm-up for both the encoder and decoder boosts accuracy by 0.6 to 1 DSC points. (ii) Learning rate adjustments matter:Reducing the peak learning rate to 1e-3 during fine-tuning consistently yields better results than the default 1e-2, with the best performance seen when fine-tuning both the encoder and decoder with lower learning rates. (iii) Freezing encoder weights is detrimental:The encoder should not remain fixed during fine-tuning. Allowing the encoder to be fine-tuned improves the model’s performance compared to when only the decoder is fine-tuned.
Transfer 1. Warm-Up 2. Warm-Up Tr. Stage Max. LR D1 D2 D3 D4 D5 Avg
[Uncaptioned image] - - [Uncaptioned image] 1e-2 45.56 72.26 88.80 60.44 82.61 69.93
[Uncaptioned image] - - [Uncaptioned image] 1e-2 49.37 69.13 88.78 60.74 81.33 69.87
[Uncaptioned image] - - [Uncaptioned image] 1e-2 50.37 70.64 88.61 61.51 81.91 70.61
[Uncaptioned image] - - [Uncaptioned image] 1e-3 49.98 71.04 88.68 61.45 82.12 70.65
[Uncaptioned image] - [Uncaptioned image] [Uncaptioned image] 1e-2 49.84 72.56 88.45 62.16 81.75 70.95
[Uncaptioned image] - [Uncaptioned image] [Uncaptioned image] 1e-3 51.54 72.74 88.85 62.44 82.33 71.58
[Uncaptioned image] - [Uncaptioned image] [Uncaptioned image] 1e-4 50.66 72.98 88.68 62.73 82.09 71.43
[Uncaptioned image] - - [Uncaptioned image] 1e-2 50.81 68.55 88.52 60.73 80.77 69.87
[Uncaptioned image] - - [Uncaptioned image] 1e-3 50.11 73.90 88.58 61.82 81.48 71.18
[Uncaptioned image] - [Uncaptioned image] [Uncaptioned image] 1e-2 51.39 72.06 88.65 61.70 81.78 71.12
[Uncaptioned image] - [Uncaptioned image] [Uncaptioned image] 1e-3 51.04 72.72 88.91 62.84 81.19 71.34
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 1e-2 48.81 72.92 88.84 62.46 82.04 71.02
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 1e-3 51.42 72.84 89.09 63.30 82.15 71.76
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 1e-4 50.13 72.26 88.72 61.70 81.93 70.95
Table 2: Fine-tuning matters. We compare various combinations of weight transfer and fine-tuning schedules. Transfer: [Uncaptioned image] Transfer all, [Uncaptioned image] Transfer encoder only. Warm-Up and Training: [Uncaptioned image] Only decoder weights trained, [Uncaptioned image] Encoder and decoder weights trained. $\ddagger$: nnU-Net default (Dynamic)
4 Results and Discussion
We compare our final models S3D-B and S3D-L against VoCo (Wu et al., 2024), VolumeFusion (VF) by Wang et al. (2023), as well as MG Models Genesis (Zhou et al., 2021). The baselines are pre-trained using the same framework on the same data with the same backbone and the same hyperparameters - where possible - and are scaled to fully utilize an A100 40GB GPU, optimized for 250k steps. We provide explicit baseline method and configuration details in Appendix A. Moreover, we compare against two from-scratch baselines. The first, ’No (Dyn.)’, represents a non-pretrained (i.e., from scratch) default nnU-Net 3D fullres architecture that was planned and preprocessed on each downstream dataset individually, potentially resulting in different architectures, data preprocessing, and spacings. The second from-scratch baseline, ’No (Fixed)’, is an nnU-Net training with the same plans and preprocessing as defined by our pre-training. The dataset-wise mean DSC and mean NSD values across our test dataset suite are provided in Table 3. Additionally, ranking stability of the methods is evaluated through bootstrapping and is provided in Fig. 3.
4.1 Observations
SSL pre-training works
Across all tested datasets, SSL pre-trained methods demonstrate improved downstream segmentation performance. Comparing our S3D-B method to the most similar from-scratch baseline ’No (Fixed)’, we observe higher DSC scores in 10 out of 11 test datasets, with an average increase of +2 DSC points and +1.6 NSD points. This improvement is not limited to our method; MG and VF also achieve higher performance than the baseline, indicating the utility of SSL methods when applied to sufficient data and a state-of-the-art architecture.
MAEs dominate
Throughout our test dataset pool, SSL schemes using the masked image modeling paradigm (MG, S3D-B, and S3D-L) consistently rank higher than the contrastive VoCo or the pseudo-segmentation-based VolumeFusion pre-training method for CNN pre-training. Given the age of ’Models Genesis’ - published in 2019 - it is surprising to see it outperform the more recent VoCo or VF. We attribute this to a combination of two factors: 1. Models Genesis was originally published and trained on an outdated 3D-UNet (Çiçek et al., 2016) and outside of the powerful nnU-Net framework (Isensee et al., 2021). This highlights the importance of avoiding Pitfall 2: Training on a state-of-the-art backbone. 2. VoCo and VF were introduced in conjunction with architectures they were optimized for. By transferring them to a CNN setting, hyperparameters chosen to optimize the method for their original architecture-pretraining combination may be suboptimal for the new CNN backbone.
S3D-B ranks first with respect to DSC and Normalized Surface Distance (NSD), while S3D-L and MG share the second place. Although S3D-L and MG are very close in Average DSC, S3D-L consistently achieves lower ranks across all datasets. Moreover, according to the bootstrapped aggregated rank (Fig. 3), this superior ranking is reliably higher than that of MG. These results indicate the overall efficacy of our pre-training method compared to currently established methods for CNN pre-training.
SSL Method No (Dyn.) No (Fixed) VoCo VF MG S3D-B S3D-L
Dataset Dice Similarity Coefficient (DSC)
MS FLAIR (D1) 57.81 59.82 59.70 59.29 58.64 60.35 59.85
Brain Mets (D2) 63.66 56.53 56.25 61.01 65.39 65.24 64.81
Hippocampus (D3) 89.18 89.24 88.78 89.03 89.38 89.60 89.34
Atlas22 (D4) 63.28 65.52 62.97 65.76 65.93 66.95 64.58
CrossModa (D5) 85.64 83.44 83.07 84.24 83.91 84.08 84.02
Cosmos22 (D6) 60.28 78.17 77.40 80.09 79.67 80.00 80.01
ISLES22 (D7) 77.94 79.44 78.14 78.96 78.85 79.70 79.89
Hanseg (D8) 59.00 61.85 57.47 61.49 62.52 62.11 61.93
HNTS-MRG24 (D9) 66.73 65.90 67.65 63.34 68.00 68.62 67.94
BRATS24 Africa (D10) 93.07 92.51 91.97 92.16 92.36 92.19 92.90
T2 Aneurysms (D11) 46.76 41.97 40.16 44.96 45.48 47.26 44.15
Avg. DSC 69.40 70.40 69.41 70.94 71.83 72.37 71.77
Avg. Rank 4.64 4.55 6.27 4.36 3.18 2.00 3.00
Dataset Normalized Surface Distance (NSD)
MS FLAIR (D1) 78.77 80.16 79.70 79.57 79.16 80.03 80.40
Brain Mets (D2) 80.72 76.72 72.77 79.20 81.51 82.53 82.32
Hippocampus (D3) 99.46 99.42 99.43 99.46 99.39 99.46 99.44
Atlas22 (D4) 70.52 73.77 70.15 73.67 74.22 75.35 73.45
CrossModa (D5) 99.85 99.76 99.72 99.78 99.74 99.81 99.80
Cosmos22 (D6) 72.60 96.47 94.48 96.89 96.95 97.45 96.75
ISLES22 (D7) 88.55 90.45 89.39 90.28 89.59 90.59 90.72
Hanseg (D8) 82.20 85.94 80.44 85.29 85.94 85.80 86.20
HNTS-MRG24 (D9) 71.83 71.26 73.47 67.88 73.22 74.07 73.17
BRATS24 Africa (D10) 95.66 95.36 94.95 94.94 95.33 95.06 95.72
T2 Aneurysms (D11) 62.24 55.56 51.79 58.97 59.38 61.18 57.07
Avg. NSD 82.04 84.08 82.39 84.17 84.95 85.58 85.00
Avg. NSD Rank 4.27 4.27 5.82 4.64 4.00 2.18 2.82
Table 3: S3D out-performs all baselines: Mean DSC and NSD results across the test datasets, representing a broad selection of brain MR tasks, are presented. ’No Fixed’ represents a from-scratch baseline sharing the same architecture, preprocessing and downstream training steps as all SSL methods. ’No Dyn.’ represents the original nnU-Net adapted to each downstream dataset individually.
Figure 3: Dataset rank stability In addition to absolute mean performance, we report the ranking stability of the methods through bootstrapping for all test datasets as well as the aggregated rank across all datasets.
Impact of dynamic configuration
Comparing the ’No (Dyn.)’ and ’No (Fixed)’ configurations, both trained from scratch, reveals that selecting the appropriate configuration for each dataset can significantly influence performance. For instance, on datasets D2 and D11, the dynamic configuration outperforms the fixed by +7 and +5 DSC points, respectively, while for D6, the fixed configuration yields results +18 DSC points higher. In the majority of datasets where the fixed configuration underperforms relative to the dynamic nnU-Net, pre-training helps to recover performance. However, in some cases, such as with D5, the dynamic default nnU-Net still proves superior.
4.2 Ablation experiments
Low-Data Regime
While previous experiments focused on comparing SSL pre-training against a from-scratch baseline with full-scale datasets available, many applications in the medical domain have access to only a very small amount of labeled images. To measure the benefits of pre-training in such a low-data regime, we artificially reduced the total amount of data available for training to 10, 20, 30, or 40 labeled images.
Results are presented in Table 4. It is evident that our pre-trained S3D-B model leads to better downstream performance compared to training from scratch in this setting. With just 40 trained images, the fine-tuned model nearly matches the performance of the from-scratch model trained on the full dataset. Future research could explore whether optimizing the training duration or learning rate schedule could prevent the pre-trained network from overfitting during fine-tuning on such a limited number of training images.
Generalization Performance
To assess the generalization capability of the proposed pre-training method, we tested two scenarios. First, we evaluated fine-tuning our method on an unseen modality using the TOF Angiography Aneurysms dataset (D12). As shown in Table 7, without pretraining, the fixed configuration suffers a performance drop of 20 Dice points. We attribute this to the significant difference in median spacing for the downstream task ($[0.50,0.43,0.43]$ mm), which has a higher resolution than the fixed target spacing of $[1,1,1]$ mm used in the pre-training experiments. This lower resolution likely increases the difficulty of segmenting small aneurysms. Despite this decrease, pre-training mitigates some of this degradation and proves highly beneficial compared to training the same configuration from scratch. Interestingly, ModelGenesis achieves the best results, potentially due to its use of intensity augmentations during pre-training, which increases robustness against brightness shifts, such as when generalizing to different MRI sequences. Second, we fine-tuned on the D2 dataset using only the T1 contrast-enhanced (T1ce) sequence and applied these models directly to D13 without any additional fine-tuning. While the dynamic configuration performed best on the in-distribution validation cases of D2, the results on D13 indicate that MAE pre-training improves generalization across different centers, with our S3D model yielding the best performance.
Pre-training Time
Previous studies have demonstrated the positive impact of extended training schedules on the quality of learned representations for downstream tasks (He et al., 2022; Feichtenhofer et al., 2022). To explore the relevance of this factor in the 3D medical domain, we conducted a similar experiment, evaluating training durations ranging from 62.5k to 1M steps. Our results indicate that the benefits of longer training schedules begin to degrade after 250k steps, as evident in Table 5. This could explain why our scaled model did not achieve further performance improvements.
Fine-tuning length
Initializing from pre-trained weights has the potential to reduce the computational resources needed for the network to adapt to new tasks. To assess this, we tested different pre-training durations on our development datasets. While maintaining 12.5k iterations for both the decoder and full network warm-up phases, we experimented with varying subsequent training lengths. As shown in Table 8, adding just 12.5k additional iterations (37.5k total) already outperforms training from scratch. However, achieving optimal performance still requires completing the full fine-tuning schedule.
Multi-Channel Input
In many medical examinations, it is common to perform multiple scans, as clinicians often require images with different characteristics for accurate decision-making. Consequently, some datasets, like D2, D8, and D10, contain multiple input modalities. While pre-training may involve all modalities, we only feed one modality at a time into the network since not all patients have scans in every modality. This raises the question of how to handle datasets with multiple registered images. To address this, we conducted a 5-fold cross-validation on the D2 development dataset. We evaluated the replication of each input modality along with random initialization of the input stem weights. Additionally, we tested freezing the stem weights during the decoder warm-up phase. As shown in Table 6, the most stable and consistently effective approach was replicating the pre-trained stem and keeping it frozen during the decoder’s warm-up period.
5 Conclusion
This work is the first to demonstrate the potential of properly configured MAEs in 3D medical image segmentation. By overcoming key pitfalls in previous research, such as limited dataset sizes, outdated architectures, and insufficient evaluation, we show a consistent performance improvement over previous SSL methods. Notably, for the first time, we achieve consistent gains over the dynamic, dataset-adaptive nnU-Net baseline, validated across a large and diverse set of development and testing datasets. While our findings are promising, several avenues remain open for future exploration. Notably, increasing training time and batch size did not lead to performance gains, but the question remains of whether scaling the pre-training dataset size or the model parameters could unlock new potential. Furthermore, the intensity shifts employed by ModelGenesis SSL task hint at intriguing possibilities for improving generalization across unseen MRI modalities, which needs to be further explored for MAEs. Lastly, a data-centric approach to curating the most relevant data for SSL represents an exciting frontier for future research. Raw clinical datasets often contain images not intended for diagnostic purposes, such as those used for scanner calibration, which can dilute the effectiveness of pre-training. While we applied basic filtering to exclude low-quality data, more sophisticated filtering techniques could significantly enhance the quality of the pre-training process.
This work follows the spirit of prior studies like nnU-Net Isensee et al. (2021), or XXX, showing that a robust development strategy, informed model configuration, and rigorous validation lead to true and sustainable performance gains, contrasting the current hype for employing and modifying the latest network architecture. With our dynamic open-source framework, we hope to contribute to a cultural shift in the community towards validation-driven development enabling true scientific progress.
Appendix A Method configuration
Across all baseline methods, we utilize a common set of hyperparameters. For all baseline methods we utilize the same pre-training dataset with the same image preprocessing. Moreover we use the same amount of pre-training steps (250k) as for our S3D-B method and the same fine-tuning scheme, as highlighted in Table 8. Aside from this, we employ the SGD optimizer with LR 1e-2 with a PolyLR schedule, momentum 0.99 and weight decay 3e-5 across all pre-training experiments, as they showed to be highly robust and reliable in the supervised medical image segmentation setting using CNNs (Isensee et al., 2021). Moreover, we denote that all these methods have their backbones replaced with a ResEnc U-Net to minimize confounding effects of different architectures.
A.1 Models Genesis
Models genesis (Zhou et al., 2021) pre-text task is centered around restoring original patches from transformed versions. The transformed version is achieved by applying four different transformations in various combinations, with the following transformations: a composition of four separate pre-training schemes: (i) Non-linear intensity transformation:Alters the intensity distribution while preserving the anatomy, focusing on learning the appearance of organs. (ii) Out-painting:Removes part of the image and requires the model to extrapolate from the remaining image, forcing it to learn the global structure of the organs. (iii) In-painting:Masks a part of the image, and the model learns to restore the missing parts, focusing on local continuity and context. After having transformed the original image through these augmentations, the model is trained to recover the original image through a convolutional encoder-decoder architecture. This approach consolidates different tasks (appearance, texture, and context learning) into one unified image restoration task, making the model more robust and generalizable.
Model specific Hyperparameters:
The entire set of hyperparameters of Models Genesis are contained within the data-augmentation. This allows us to transfer this transformation pipeline, as provided in the official https://github.com/MrGiovanni/ModelsGenesis without any changes to the hyperparameters.
A.2 VolumeFusion
Volume Fusion (Wang et al., 2023) is a pseudo-segmentation task using two sub-volumes from different 3D scans, which are fused together based on random voxel-level fusion coefficients. The fused image is treated as input, and the model predicts the fusion category of each voxel, mimicking a segmentation task. Pretraining is optimized using a combination of Dice loss and cross-entropy loss.
Method specific parameters:
Volume Fusion has unique parameters defining the size ranges of the rectangles used for fusing together images. In our experiments we utilize a rectangle size range between [8, 100] sampled uniformly for each axis. This represent the 62.5% of our input patch size, and identical percentage as in the original paper. Moreover the amount of rectangles sampled is an important parameter. Like in the original paper we sample M Ũ(10, 40) different rectangles, iteratively. Lastly, the number of categories was chosen to be 5, as in the original paper (this represent K = 4).
A.3 VoCo
The ’Volume Contrastive Learning Framework’ (VoCo) (Wu et al., 2024) is designed to enhance self-supervised learning for 3D medical image analysis by leveraging the consistent contextual positions of anatomical structures. The method involves generating base crops from different regions of 3D images and using these as class assignments. The framework then contrasts random sub-volume crops against these base crops, predicting their contextual positions using a contrastive learning approach. The authors utilize a Swin-UNETR model architecture, employing the AdamW optimizer with a cosine learning rate schedule for 100,000 pre-training steps. The specific hyperparameters include cropping non-overlapping volumes with a size of 64x64x64, and generating 4x4 base crops during the position prediction task. This represents an input patch size of 384×384×96 which is rescaled and resized to fit exactly 4x4 64x64x64 crops.
Since our chosen patch size 160x160x160 is incompatible with the 64 cube length, we adjusted our patch size for VoCo to 192x192x64. This accommodates a 3x3 grid of 64x64x64. Unfortunately the 4x4 grid led to exceeding the memory limit hence a reduction was necessary. Moreover we increased the target crop size from 4 originally to 5 and increased the batch size from 6 (default in our other experiments) to 12, to fully utilize the 40GB VRAM of an A100 node.
Appendix B Additional results
Aside from the quantitative data on the development and test dataset, we provide the quantiative data of the ablation experiments here. The following additional results are provided: 1. Results when fine-tuning in a low-data regime are presented in Table 4. 2. Experiment on how to best transfer weights when transferring to a dataset with more than 1 input channel is provided in Table 6 3. Results on how the pre-training effects generalization is provided in Table 7. 4. Experiment results of investigating if one can reduce the fine-tuning steps are presented in Table 8
SSL Method N Train D1 D2 D3 D4 D5 Avg. D1-D5
Scratch 10 40.78 43.52 84.94 44.11 76.66 58.00
20 44.46 59.46 86.75 46.33 78.67 63.13
30 45.42 64.20 87.14 48.22 78.47 64.69
40 49.37* 60.13 87.59 50.43 78.37 65.18
full 49.37 69.13 88.78 60.74 81.33 69.87
S3D-B (ours) 10 43.48 48.44 84.12 41.51 77.70 59.05
20 46.58 65.30 86.61 45.50 79.52 64.70
30 48.12 68.41 86.77 51.62 78.88 66.76
40 51.49* 72.91 87.46 53.05 80.82 69.15
full 51.49 74.01 88.83 62.39 81.54 71.65
Table 4: Forty images with SSL are almost as good as all data from-scratch! S3D-B model almost reaches the performance of the model trained from-scratch with only 40 training cases. * D1 has only 38 training cases for the train split.
Table 5: Pre-training length ablation: Longer pre-training does not lead to improved performance. Interestingly, when exceeding 250k steps.
PT Iterations D1 D2 D3 D4 D5 Avg. D1-D5 Train Time [h]
62.5k 49.49 70.79 88.82 62.95 81.27 70.67 28
125k 50.56 70.48 88.86 62.51 81.69 70.82 56
250k 51.02 74.07 88.91 62.81 81.50 71.66 112
500k 50.93 72.71 88.88 62.17 81.86 71.31 224
1M 50.45 71.55 88.92 62.78 81.82 71.10 448
Initialization Decoder Warm-Up Fold 0 Fold 1 Fold 2 Fold 3 Fold 4 Average STD
Replication Frozen 72.84 64.42 66.11 62.86 62.85 65.82 4.15
Replication Unfrozen 72.68 63.07 65.60 66.02 61.08 65.69 4.39
Random Frozen 74.38 60.89 65.10 67.55 61.31 65.85 5.51
Random Unfrozen 72.20 63.16 62.25 66.71 61.47 65.16 4.42
Table 6: Replicating the pre-trained stem weights and freezing them during the decoder warm-up phase yields the most stable and equally best results.
Experiment Setting No Dyn. No Fixed VoCo VF MG S3D-B S3D-L
Modality shift TOF Angio. Aneurysms(D12) 42.61 22.76 22.32 31.21 34.60 28.72 26.90
In Distribution Brain Mets (D2) 72.81 67.93 64.34 71.69 69.05 71.56 72.53
Patient shift Brain Mets (D13) 64.08 61.61 56.78 63.95 64.22 64.54 64.95
Table 7: Pre-training can improve generalization: We investigate generalization to a new modality time-of-flight (ToF) MRI (top), and the generalization of a resulting method when translating it to a different clinic (bottom).
Table 8: Fine-tuning length: When initializing from our pre-trained checkpoint, it is possible to achieve a large fraction of the final performance after less than 15% of the normal training time. Despite this a full training schedule reaches better performance.
FT Iterations D1 D2 D3 D4 D5 Avg. D1-D5
25k 50.85 73.99 88.51 55.49 46.00 62.97
37.5k 51.69 74.03 88.85 60.22 81.68 71.29
50k 51.13 73.53 88.93 60.14 81.92 71.13
75k 51.41 72.80 89.08 63.14 81.83 71.65
150k 50.95 71.28 88.96 62.51 81.92 71.13
275k 53.10 71.24 89.14 63.55 82.53 71.91

### v2

@@
 Self-Supervised Learning (SSL) presents an exciting opportunity to unlock the potential of vast, untapped clinical datasets, for various downstream applications that suffer from the scarcity of labeled data.
-While SSL has revolutionized fields like natural language processing and computer vision, their adoption in 3D medical image computing has been limited by three key pitfalls: Small pre-training dataset sizes, architectures inadequate for 3D medical image analysis, and insufficient evaluation practices.
-We address these issues by i) leveraging a large-scale dataset of 44k 3D brain MRI volumes and ii) using a Residual Encoder U-Net architecture within the state-of-the-art nnU-Net framework.
+While SSL has revolutionized fields like natural language processing and computer vision, its adoption in 3D medical image computing has been limited by three key pitfalls: Small pre-training dataset sizes, architectures inadequate for 3D medical image analysis, and insufficient evaluation practices.
+In this paper, we address these issues by i) leveraging a large-scale dataset of 39k 3D brain MRI volumes and ii) using a Residual Encoder U-Net architecture within the state-of-the-art nnU-Net framework.
 iii) A robust development framework, incorporating 5 development and 8 testing brain MRI segmentation datasets, allowed performance-driven design decisions to optimize the simple concept of Masked Auto Encoders (MAEs) for 3D CNNs.
-The resulting model not only surpasses previous SSL methods but also outperforms the strong nnU-Net baseline by an average of approximately 3 Dice points.
-Furthermore, our model demonstrates exceptional stability, achieving the highest average rank of 2 out of 7 methods, compared to the second-best method’s mean rank of 3.
-Our code is made available here.
+The resulting model not only surpasses previous SSL methods but also outperforms the strong nnU-Net baseline by an average of approximately 3 Dice points setting a new state-of-the-art.
+Our code and models are made available here.
 1 Introduction
-In recent years, the concept of Self-Supervised Learning (SSL) has emerged as a driving factor in data-rich domains, enabling large-scale pre-training that facilitates the learning of robust and transferable general-purpose representations (Assran et al., 2023; Oquab et al., 2023; He et al., 2022).
+Figure 1: Well-configured MAE pre-training for CNNs is state-of-the-art.
+When comparing our MAE pre-trained model on all eleven test Datasets, Spark3D achieves almost +3 DSC points compared to the strong nnU-Net baseline and outperforms current SSL methods.
+In recent years, the concept of Self-Supervised Learning (SSL) has emerged as a driving factor in data-rich domains, enabling large-scale pre-training that facilitates the learning of robust and transferable general-purpose representations [2, 25, 14].
 This paradigm shift has been instrumental in advancing various fields, particularly in domains with abundant labeled data like NLP or natural vision.
@@
 Currently, the domain is either focused on training from-scratch, mainly using the nnU-Net framework by Isensee et al.
-(2021), or using supervised pretraining, which is limited by the cost associated with annotated data (Wasserthal et al., 2023; Ulrich et al., 2023; Huang et al., 2023).
-The usage of supervised pretraining implies the willingness of the domain to adopt pretraining and calls to question the currently established SSL methods in the domain.
+[18], or using supervised pre-training, which is limited by the cost associated with annotated data [40, 35, 17].
+The usage of supervised pre-training implies the willingness of the domain to adopt pre-training and calls to question the currently established SSL methods in the domain.
 We believe this lack of widespread adoption of previously established methods can be attributed to three major pitfalls of previous SSL research in this field:
-P1 - Limited dataset size: Many SSL approaches have been developed on very few unlabeled volumes, often being trained on fewer than 10,000 images Zhou et al.
-(2021); Wu et al.
-(2024); Tang et al.
-(2024; 2022); Zhuang et al.
-(2023); Hatamizadeh et al.
-(2021), almost approaching scales of supervised dataset sizes.
+P1 - Limited dataset size: Many SSL approaches have been developed on very few unlabeled volumes, often being trained on fewer than 10,000 images [44, 42, 30, 31, 45, 12], almost approaching scales of supervised dataset sizes.
 These datasets tend to be pooled from publicly available annotated datasets, as larger datasets pose a greater hurdle to acquire.
 While many hospitals possess 3D medical images in the millions, they are locked away from the public due to patient privacy concerns.
-While some larger open-source datasets exist, e.g., the Adolescent Brain Cognitive Decline (ABCD) dataset of the NIH (N=40k) or the UK Biobank (UK-BB) (N=120k), they restrict access pending an internal review board’s approval, posing a hurdle for open science.
-More recently, the UK-BB stopped allowing downloading data to local hardware, signaling a public decrease in the community’s willingness to share data.
-P2 - Outdated Backbones: Many studies develop their SSL method on non-state-of-the-art architectures, e.g., utilizing transformers (Wu et al., 2024; Tang et al., 2022; Wang et al., 2023; Chen et al., 2023).
-While transformers are prevalent in the 2D natural imaging domain (Dosovitskiy, 2020), recent architectures leveraging attention (Vaswani, 2017) have so far not been able to reach state-of-the-art performance in 3D medical segmentation.
-In fact, well-configured 3D U-Net (Çiçek et al., 2016; Ronneberger et al., 2015) inspired CNNs dominate 3D medical image segmentation, outperforming transformer-based models by a large margin Isensee et al.
-(2024).
+Some larger open-source datasets exist, e.g., the Adolescent Brain Cognitive Decline (ABCD) dataset of the NIH (N=40k) or the UK Biobank (UK-BB) (N=120k), but they restrict access pending an internal review board’s approval, posing a hurdle for open science.
+P2 - Outdated Backbones: Many studies develop their SSL method on non-state-of-the-art architectures, e.g., utilizing transformers [42, 31, 39, 5].
+While transformers are prevalent in the 2D natural imaging domain [8], recent architectures leveraging attention [36] have so far not been able to reach state-of-the-art performance in 3D medical segmentation.
+In fact, well-configured 3D U-Net [6, 27] inspired CNNs dominate 3D medical image segmentation, outperforming transformer-based models by a large margin [19].
 This underscores the need for SSL methods that can be seamlessly integrated with CNNs to harness their full potential in medical image analysis on downstream tasks.
-P3 - Insufficient Evaluation: Existing methods often lack rigorous evaluation, masking the methods’ efficacy (or lack thereof).
-This is represented through: i) Evaluating on too few datasets to show generalization of the pre-training ii) Stacking multiple contributions, e.g., novel architecture with new pre-training, which does not allow one to draw conclusions if the pre-training is effective without the architecture (Wang et al., 2023) iii) Comparing to bad from-scratch baselines, like a badly configured, outdated Çiçek et al.
-(2016) model instead of a well-configured 3D nnU-Net CNN baseline (Isensee et al., 2021).
+P3 - Insufficient Evaluation: Existing methods often lack rigorous evaluation, masking the methods’ efficacy.
+This is represented through: i) Evaluating on too few datasets to show generalization of the pre-training ii) Stacking multiple contributions, e.g., novel architecture with new pre-training, which does not allow one to draw conclusions if the pre-training is effective without the architecture [39] iii) Comparing to bad from-scratch baselines, like a badly configured, outdated [6] model instead of a well-configured 3D nnU-Net CNN baseline [18].
 iv) Evaluating their method on seen data they pre-trained on.
@@
 This observation is similar to the recent study by Isensee et al.
-(2024), that shows that this is also prevalent in the field of medical image segmentation when training from-scratch.
-In this paper, we carefully avoided all these pitfalls while exploring the Masked Auto Encoder (MAE) paradigm for 3D CNNs, with the recent adaptations introduced by Tian et al.
-(2023); Woo et al.
-(2023), and highlight that MAEs are exceeding the current state-of-the-art SSL methods in 3D medical image segmentation given proper configuration.
-Our contributions can be summarized as follows:
+[19], that shows that this is also prevalent in the field of medical image segmentation when training from-scratch.
+Table 1: Current SSL methods are limited by three pitfalls: They are trained on less than 10K 3D images (P1), they do not use a state-of-the-art backbone for 3D image segmentation as recommended by Isensee et al.
+[19] (P2) or they do not evaluate their method on more than 4 downstream datasets (P3).
+[Uncaptioned image] indicates evading a pitfall while [Uncaptioned image] indicates falling for a pitfall.
+Method Year P1 P2 P3
+Data Model Eval.
+3DMAE [5] Jan. 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+GL-MAE [45] Aug. 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+MAPSeg [43] March 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+HySparK [30] Aug. 2024 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+AMAES [23] Aug. 2024 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Swin UNETR [31] March 2022 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+VoCo [42] April 2024 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+VF [39] June 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+MG [44] Aug. 2019 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+GVSL [15] March 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+S3D (ours) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+In this paper, we explore the Masked Auto Encoder (MAE) paradigm for 3D CNNs, with the recent adaptations introduced by Tian et al.
+[32], Woo et al.
+[41].
+Our approach addresses the key pitfalls of previous SSL research in 3D medical image segmentation, ensuring broader applicability and higher performance.
+Our main contributions are as follows:
   1. 1.
-    We evade Pitfall 1 by leveraging a dataset collection of 44k 3D MRI volumes to develop our self-supervised pretraining, exceeding the majority of 3D medical image segmentation SSL methods in scale (Zhou et al., 2021; Wu et al., 2024; Tang et al., 2024; 2022; Zhuang et al., 2023; Hatamizadeh et al., 2021).
+    State-of-the-Art Performance: Our final model achieves superior performance compared to existing SSL-pre-trained models using the same architecture and dataset, see Fig.
+1.
+Notably, this is the first work to demonstrate that SSL pretraining with a fixed architecture can consistently outperform a state-of-the-art, dynamically optimized nnU-Net baseline by approximately 3 Dice points across 11 diverse downstream datasets [18].
+An associated model will be made available.
   2. 2.
-    We evade Pitfall 2 by utilizing the state-of-the-art Residual Encoder U-Net CNN architecture from Isensee et al.
-(2024) as the backbone.
-Moreover, we use this backbone for all baseline SSL methods - allowing us to quantify the utility of the pre-training scheme for pre-training CNNs.
+    Comprehensive Evaluation Addressing Previous Pitfalls: Building on the detailed analysis of limitations in prior work (see Tab.
+1), we set up a robust development environment to design our method.
+() P1: We utilize a collection of 39k 3D MRI volumes for self-supervised pre-training, substantially larger than most datasets used in prior SSL work [44, 42, 30, 31, 45, 12].
+() P2: By using the state-of-the-art Residual Encoder U-Net CNN architecture [19], we bypass the limitations of outdated or transformer-based architectures, leveraging CNNs’ superior performance in 3D medical segmentation.
+() P3: We rigorously evaluate our model across five development and eight testing datasets, covering a wide range of downstream tasks, diverse pathologies, novel image modalities, and multi-center datasets.
+Moreover, we disentangle pre-training methodology from other contributions like novel backbones and include the strong 3D nnU-Net baseline, ensuring reliable results.
   3. 3.
-    We evade Pitfall 3 by employing 5 development and 7 testing datasets, spanning a diverse set of downstream targets.
-This includes head and neck organs and pathologies, datasets with novel image modalities not seen during pre-training, and datasets of the same pathology acquired at different centers.
-By subsequently optimizing MAEs for CNNs based on performance on the development pool, we propose the novel sparse MAE inspired pre-training paradigm, unleashing the potential of huge amounts of not annotated data in the medical domain.
-We show that the choice of the best fine-tuning strategy is crucial.
-Moreover, we evaluate our method across a range of critical 3D medical scenarios, including low-data regimes, accelerated fine-tuning schedules, and generalization across unseen centers and modalities.
-Figure 1: Overview of the model development pipeline.
-During pretraining, we carefully develop our method with 44k images (P1), with a state-of-the-art residual encoder U-Net (P2) and evaluate it on dive development datasets (P3).
-In the testing phase we utilize a total of thirteen datasets (P3) to verify the utility of our developed method.
+    In-Depth Ablation Studies: We provide detailed analyses of critical design decisions, such as masking ratio, sparsification, and fine-tuning strategies.
+Additional ablations explore generalization, low-data performance, and faster fine-tuning, offering valuable insights.
+While the fundamental concept of an MAE is not novel, its application in 3D medical image analysis has yet to be fully validated due to limitations in previous studies, which we illustrate in Tab.
+1.
+Our carefully designed experiments close these gaps and reveal MAEs true potential.
 2 Development Framework
@@
 To reduce the search space and disentangle the effects of SSL pre-training from other basic design choices, we choose to keep some parameters fixed based on best practices in the domain:
-(i) The architecture used is always the same state-of-the-art residual encoder U-Net architecture (ResEnc U-Net) (Isensee et al., 2024).
-(ii) The input patch size is [160x160x160].
-(iii) All images are resampled to the target spacing of [1x1x1] mm3 (Roy et al., 2023).
-(iv) All images are z-score normalized to zero mean and unit variance (Isensee et al., 2021).
-(v) As optimizer, we use SGD with a decreasing poly-learning rate (Chen et al., 2018) following nnU-Net (Isensee et al., 2021).
+(i) The architecture used is always the same state-of-the-art residual encoder U-Net architecture (ResEnc U-Net) [19, 35].
+(ii) The input patch size is set to [160x160x160].
+(iii) All images are resampled to the same target spacing of [1x1x1] mm [28].
+(iv) All images are z-score normalized to zero mean and unit variance [18].
+(v) As optimizer, we use SGD with a decreasing poly-learning rate [4] following nnU-Net [18].
 (vi) We always employ random sampling during pre-training, irrespective of the prevalence of the different MR modalities.
@@
 Due to the variety of data sources, this dataset contains images from more than 10 different MR scanners, various MR modalities and a diverse patient population.
-For more details on the distributions we refer to Figure 2.
+For more details on the distributions, we refer to Figure 3.
 Since this data is sourced directly from clinical examinations, it includes empty or broken images, poor-quality images and so-called scout scans used to determine the field of view of the patient in the MR.
@@
   1. 1.
-    MS FLAIR (D1): Consensus delineations of multiple sclerosis (MS) lesions on T2-weighted FLAIR images(Muslim et al., 2022).
+    MS FLAIR (D1): Consensus delineations of multiple sclerosis (MS) lesions on T2-weighted FLAIR images [24].
   2. 2.
-    Brain Mets (D2): Brain metastases imaged through T1, contrast enhanced gradient echo T1ce, contrast enhanced spin echo T1 and a T2 FLAIR sequence acquired in the Stanford University Hospital (Grøvik et al., 2020).
+    Brain Mets (D2): Brain metastases imaged through T1, contrast enhanced gradient echo T1ce, contrast enhanced spin echo T1 and a T2 FLAIR sequence acquired in the Stanford University Hospital [11].
   3. 3.
-    Hippocampus (D3): The Hippocampus dataset, task 4 of the medical segmentation decathlon (MSD) (Antonelli et al., 2022), contains delineations of the anterior and posterior Hippocampus in T1 weighted MRI (Simpson et al., 2019).
+    Hippocampus (D3): The Hippocampus dataset, task 4 of the medical segmentation decathlon (MSD) [1], contains delineations of the anterior and posterior Hippocampus in T1 weighted MRI [29].
   4. 4.
-    Atlas22 (D4): Anatomical Tracings of Lesions After Stroke (ATLAS) on T1 weighted images. We use the Atlas R2.0 dataset from Liew et al. (2022).
+    Atlas22 (D4): Anatomical Tracings of Lesions After Stroke (ATLAS) on T1 weighted images. We use the Atlas R2.0 dataset from Liew et al. [22].
   5. 5.
-    CrossModa (D5): Delineations of intra-meatal and extra-meatal vestibular schwannoma tumors and cochlea delineations in contrast enhanced T1 weighted MRI (Dorent et al., 2023).
-Of all these datasets, we set aside a hold-out test set comprising 20% of all images before start of the method development process.
-The remaining images were further split 80/20 into training and validation sets for the development process.
+    CrossModa (D5): Delineations of intra-meatal and extra-meatal vestibular schwannoma tumors and cochlea delineations in contrast-enhanced T1 weighted MRI [7].
+Of all these datasets, we set aside a hold-out test set comprising 20% of all images before the start of the method development process.
+The remaining images were split 80/20 into training and validation sets for the development process.
 Test Datasets
@@
   1. 1.
-    Cosmos (D6): This dataset features carotid vessel wall segmentation and atherosclerosis diagnosis, of which we use the contours to evaluate segmentation performance (Chen et al., 2022).
+    Cosmos (D6): This dataset features carotid vessel wall segmentation and atherosclerosis diagnosis, of which we use the contours to evaluate segmentation performance [3].
   2. 2.
-    HaNSeg (D7): This dataset contains segmentations of 30 organs at risk (OAR), with associated T1 MRI and CT, of which we only use the MR images for model development (Podobnik et al., 2023).
+    HaNSeg (D7): This dataset contains segmentations of 30 organs at risk (OAR), with associated T1 MRI and CT, of which we only use the MR images for model development [26].
   3. 3.
-    Isles22 (D8): This dataset contains annotations of ischemic stroke lesions, with associated diffusion-weighted imaging (DWI), apparent diffusion coefficient (ADC) and T2 FLAIR (Hernandez Petzsche et al., 2022).
+    Isles22 (D8): This dataset contains annotations of ischemic stroke lesions, with associated diffusion-weighted imaging (DWI), apparent diffusion coefficient (ADC) and T2 FLAIR [16].
   4. 4.
-    HNTS-MRG24(D9): T2-weighted MR images of pre-treatment oropharyngeal cancer and metastatic lymph nodes with associated annotations (Wahid et al., 2024).
+    HNTS-MRG24 (D9): T2-weighted MR images of pre-treatment oropharyngeal cancer and metastatic lymph nodes with associated annotations [38].
   5. 5.
@@
 However, instead of fine-tuning on it we use it to measure generalization when inferring models trained on D2 on it.
-For all test datasets that are fine-tuned on (D6-D11) we use an 80/20 split for fine-tuning and for testing, as we fine-tune each method only once without any interventions.
-For D13 all data is used, as no training is conducted.
-[t]
-Figure 2: Hierarchical composition of the nnSSL foundation model’s pre-training dataset.
-(a) Geographic distribution of 44 participating centers.
-(b) Patient demographics (age, sex, weight).
-(c) Scanner specifications (magnetic field strength, manufacturer, scanner) of examinations.
-(d) Acquisition parameters (orientation, contrast administration, weighting) of image volumes.
+For test datasets D6-D12 we use an 80/20 split for fine-tuning and for testing.
+A validation set is omitted as we fine-tune each method only once without any interventions.
+For D13 all data is used for testing, as no training is conducted.
+3 Related Work
+In this section, we revisit currently established MAE pre-training paradigms.
+Then, we survey existing MAE pre-training paradigms in the 3D medical domain, before expanding this to a broader set of pre-training paradigms for 3D medical image segmentation.
+3.1 Masked Autoencoders
+Auto-encoders have been established for a long time through the early pioneering works of Vincent et al.
+[37] but recently garnered significant attention through He et al.
+[14], showing that a simple MAE pre-training paradigm is scalable and provides substantial performance improvements.
+Subsequently, follow-up works extended these findings to video data [9, 34] as well as language-image data [21].
+These earlier successes primarily relied on Transformer architectures, as their sequence modeling offered significant VRAM savings.
+CNNs on the other hand are less suited to such a pre-training paradigm as masking disrupts the 2D data-structure they compute on.
+Despite this Woo et al.
+[41] and Tian et al.
+[32] were able to extend MAE pre-training successfully to CNNs.
+Woo et al.
+[41] introduced their fully convolutional masked autoencoder (FCMAE) framework, while Tian et al.
+[32] introduced adaptations to the convolution computation and normalization to make it resilient to masking.
+3.2 Masked Autoencoder In 3D Medical
+The earliest work using autoencoders for self-supervised pre-training in the medical domain was Models Genesis (MG) conducted by Zhou et al.
+[44] in 2019.
+They introduced reconstruction of in-painting, out-painting, and denoising of local intensity shifts as their pre-training task.
+Many approaches followed, with Swin UNETR by Tang et al.
+[31] proposing a joint in-painting, rotation, and contrastive objective for Swin Transformer pre-training.
+Chen et al.
+[5] proposed a basic masking and reconstruction strategy applied to 3D Vision Transformers (ViT) [8] termed (3DMAE) and Zhuang et al.
+[45] proposed to use a combination of masked global and local views in conjunction with a reconstruction and consistency loss to pre-train 3D ViTs, named GL-MAE.
+MAPSeg by Zhang et al.
+[43] combines a masked autoencoder objective and pseudo-labeling to address domain shifts in a federated setting.
+More recently, in concurrent work, Munk et al.
+[23] proposed pre-training CNNs with a vanilla masking and reconstruction objective without any adaptations on a large brain dataset (AMAES) and Tang et al.
+[30] presented a sparse MAE with reconstruction objective pre-trained on abdomen CT data (HySparK).
+3.3 Other Pre-Training For 3D Medical
+Additionally, self-supervised learning approaches for 3D medical images include Zhuang et al.
+[46], who introduced a training objective aimed at recovering the orientation and rotation of shuffled and rotated sub-volumes, inspired by a Rubik’s cube.
+Wang et al.
+[39] proposed a method that involves mixing two images at varying patch rates and predicting the mixing categories via a segmentation objective, termed Volume Fusion (VF).
+Geometric Visual Similarity Learning (GVSL) of He et al.
+[15] leveraged anatomical consistency by learning to register two images of the same field of view and using local similarity as a training objective.
+More recently, Wu et al.
+[42] suggested partitioning a volume into non-overlapping base-crops and generating additional overlapping sub-crops, with a training objective to predict the degree of overlap between sub-crops and base-crops (VoCo).
 4 Revisiting 3D MAEs
+Table 2: Development Experiments: During development, we evaluate the Average DSC on all development datasets to quantify the best configuration.
+Methods of the same configuration are denoted through common colors.
+(a) Sparsification: The introduction of all sparsification components showed best results, with most improvements from adding the additional densification conv layer between encoder and decoder.
 Configuration D1 D2 D3 D4 D5 Avg. D1-D5
@@
 + w/ Dens. Conv. 51.02 74.07 88.91 62.81 81.50 71.66
-(a) Sparsification: Introducing all sparsification adaptations showed best results, with most gains from adding the additional densification conv layer between encoder and decoder.
+(b) Masking ratio: Masking ratios between 60% and 75% worked best when choosing static ratios. Overall, a dynamic range, including higher masking ratios, performs the best.
 Mask ratio D1 D2 D3 D4 D5 Avg. D1-D5
@@
 U[60%-90%] 51.49 74.01 88.83 62.39 81.54 71.65
-(b) Masking ratio: Masking ratios between 60% and 75% worked best when choosing static ratios with a dynamic range including higher masking ratios performing the best.
-Pretraining D1 D2 D3 D4 D5 Avg. D1-D5
-No Dyn. 45.56 72.26 88.80 60.44 82.61 69.93
-No Fixed 49.37 69.13 88.78 60.74 81.33 69.87
-VoCo 50.35 67.20 88.22 57.82 80.29 68.77
-VF 49.93 69.58 88.83 61.75 81.48 70.31
-MG 50.50 71.14 88.83 63.29 82.15 71.18
-S3D-B (ours) 51.49 74.01 88.83 62.39 81.54 71.65
-S3D-L (ours) 51.42 72.84 89.09 63.30 82.15 71.76
-(c) Compound scaling: When comparing our scaled S3D-L, trained for more iterations with higher batch size and lr, only a modest increase in performance can be observed over the much more compute-efficient S3D-B.
-Moreover, when comparing models on the same scale S3D-B exceeds all baselines on our development validation datasets.
-Table 1: Development Experiments: During development, we evaluate the Average DSC on all development datasets to quantify the best configuration.
-Methods of the same configuration are denoted through common colors.
+(c) Development performance: When comparing our final S3D model against the baseline methods, trained equally, S3D exceeds all baselines on the development datasets.
+SSL Method No (Dyn.) No (Fix.) S3D* S3D
+Dataset Dice Similarity Coefficient (DSC)
+MS FLAIR (D1) 57.81 59.82 59.75 60.35
+Brain Mets (D2) 63.66 56.53 64.20 65.24
+Hippocampus (D3) 89.18 89.24 89.45 89.60
+Atlas22 (D4) 63.28 65.52 66.61 66.95
+CrossModa (D5) 85.64 83.44 83.61 84.08
+Cosmos22 (D6) 60.28 78.17 80.01 80.00
+ISLES22 (D7) 77.94 79.44 78.94 79.70
+Hanseg (D8) 59.00 61.85 61.27 62.11
+HNTS-MRG24 (D9) 66.73 65.90 67.03 68.62
+BRATS24 Africa (D10) 93.07 92.51 92.49 92.19
+Avg. DSC 71.66 73.24 74.34 74.88
+Avg. Rank 3.2 2.9 2.4 1.5
 Masked autoencoders (MAEs) are a well-established pre-training paradigm in the natural imaging domain and in the medical image segmentation domain for transformers.
 In this section, we investigate this paradigm and optimize it for 3D medical image segmentation using a ResEnc U-Net architecture of Isensee et al.
-(2024).
+[19].
 Default parameters
-MAEs are trained by masking an input image to a certain degree and training the network to reconstruct the occluded regions, minimizing deviations betweem reconstruction and original image.
-In our experiments, we train the MAE with a L2-Loss in the z-score normalized voxel space and only calculate the reconstruction loss where regions were masked.
+MAEs are trained by masking an input image to a certain degree and training the network to reconstruct the occluded regions, minimizing deviations between reconstruction and the original image.
+In our experiments, we train the MAE with an L2-Loss in the z-score normalized voxel space and only calculate the reconstruction loss where regions were masked.
 Moreover, we do not remove skip-connections, following the general consensus of Woo et al.
-(2023), Tian et al.
-(2023), and He et al.
-(2022).
-The default hyperparameters (as used by the model denoted in gray in Table 1) are learning rate 1e-2, weight decay 3e-5, batch size 6, SGD optimizer with Nesterov momentum 0.99, masking ratio of 75% trained with a PolyLR schedule for 250k steps (this represent 1000 epochs in the nnU-Net framework) and minor spatial augmentations of affine scaling, rotation and mirroring.
+[41], Tian et al.
+[32], and He et al.
+[14].
+The default hyperparameters (as used by the model denoted in gray in Tab.
+2(c)) are learning rate 1e-2, weight decay 3e-5, batch size 6, SGD optimizer with Nesterov momentum 0.99, masking ratio of 75% trained with a PolyLR schedule for 250k steps (this represent 1000 epochs in the nnU-Net framework) and minor spatial augmentations of affine scaling, rotation and mirroring.
 Sparsification
@@
 To address this, Tian et al.
-(2023) proposed to adapt the CNN architectures to better fit the sparse inputs: (a) Sparse Convolutions and Normalization:Through the receptive field of convolutions masked-out regions are iteratively eroded from their boundaries.
+[32] proposed to adapt the CNN architectures to better fit the sparse inputs: (a) Sparse Convolutions and Normalization:Through the receptive field of convolutions masked-out regions are iteratively eroded from their boundaries.
 By re-applying the masked regions after every convolution this problem can be resolved.
@@
 (c) Densification Convolution:After filling the masks with the Mask Token and before passing the feature maps to the decoder, a [3x3x3] convolution is applied to the feature maps at every resolution except the highest resolution to prepare the representations for decoding.
-Results of these changes are visualized in Table 1(a).
+Results of these changes are visualized in Tab.
+2(a).
 The adaptations are introduced iteratively, meaning the ’MaskToken’ ablation is only applied together with the Sparse Convolutions and Normalization.
@@
 This results in masking regions of [32x32x32] voxels of non-overlapping regions in the input.
-As a sampling strategy, we follow random masking, as previous work showed no benefit of structured masking for images nor videos (He et al., 2022; Feichtenhofer et al., 2022).
+As a sampling strategy, we follow random masking, as previous work showed no benefit of structured masking for images nor videos [14, 9].
 In the scope of the development phase we explore 5 static masking ratios between 30 and 90%, and evaluate a dynamic masking ratio randomly masking between 60% and 90%.
-Results are presented in Table Table 1(b) and highlight that the masking ratios of 60%, 75%, and the dynamic masking ratio of 60% to 90% perform equally well.
+Results are presented in Table Tab.
+2(b) and highlight that the masking ratios of 60%, 75%, and the dynamic masking ratio of 60% to 90% perform equally well.
 Due to the highly similar performance, we choose to proceed with a dynamic masking ratio over the static masking ratio, due to expecting this masking to be more difficult to learn and with the upcoming scaling experiment in mind.
-We refer to this model as Spark3D-Base (S3D-B= going forward.
-Scaling
-MAEs are known to benefit from scaling.
-We evaluate the effect of compound scaling by increasing the batch size by x8 to $48$, the learning rate to $23e-$ 2, and the iterations by x4 to $1M$.
-We refer to this model as S3D-Large (S3D-L) to denote the higher compute resources, but note that the architecture remains identical, to maintain easy adaptation of the parameters.
-Results are presented in Table 1(c).
-It can be observed that this x32 increase in compute only resulted in a slight increase in performance of 0.1 DSC points.
-Fine-tuning strategy
-Given a pre-trained model, a crucial question arises: Which weights to transfer and how to schedule the fine-tuning?
-We investigated various different schedules.
-Regarding weight transfer we investigated transferring (i) both the encoder and decoder [Uncaptioned image], or (ii) only the encoder [Uncaptioned image] with a randomly initialized decoder.
-Regarding the fine-tuning schedule, we investigate whether to use a learning rate warm-up of 12.5k steps, ramping the learning rate up to the maximum LR.
-When only transferring the encoder, an additional warm-up of only the decoder is investigated to adapt the randomly initialized decoder to the pre-trained encoder.
-In some configurations, this results in two learning rate warm-ups of 12.5k steps each.
-Additionally, we investigate whether to keep the encoder frozen [Uncaptioned image] for the entire fine-tuning process, or to fine-tuning encoder weights as well.
-Lastly, we investigate whether to decrease the peak learning rate to $31e-$ 3, $41e-$ 4 or to keep it at the default of $21e-$ 2.
-Results are presented in Table 2 and allow three important observations to be made: (i) Warm-up stages are essential:Not applying a warm-up step significantly reduces performance.
-Including a warm-up for both the encoder and decoder boosts accuracy by 0.6 to 1 DSC points.
-(ii) Learning rate adjustments matter:Reducing the peak learning rate to 1e-3 during fine-tuning consistently yields better results than the default 1e-2, with the best performance seen when fine-tuning both the encoder and decoder with lower learning rates.
-(iii) Freezing encoder weights is detrimental:The encoder should not remain fixed during fine-tuning.
-Allowing the encoder to be fine-tuned improves the model’s performance compared to when only the decoder is fine-tuned.
-Transfer 1. Warm-Up 2. Warm-Up Tr. Stage Max. LR D1 D2 D3 D4 D5 Avg
+We refer to this model as Spark3D (S3D) going forward.
+Table 3: Fine-tuning matters.
+We compare various combinations of weight transfer and fine-tuning schedules.
+Transfer: [Uncaptioned image] Transfer all weights, [Uncaptioned image] Transfer encoder weights only.
+Warm-Up and Fine-tuning: [Uncaptioned image] Only decoder weights adapted during fine-tuning, [Uncaptioned image] Encoder and decoder weights adapted during fine-tuning.
+$\ddagger$: nnU-Net default (Dynamic planning.)
+Transfer 1. Warm-Up 2. Warm-Up Fine-tuning Max. LR D1 D2 D3 D4 D5 Avg
 [Uncaptioned image] - - [Uncaptioned image] 1e-2 45.56 72.26 88.80 60.44 82.61 69.93
@@
 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 1e-4 50.13 72.26 88.72 61.70 81.93 70.95
-Table 2: Fine-tuning matters.
-We compare various combinations of weight transfer and fine-tuning schedules.
-Transfer: [Uncaptioned image] Transfer all, [Uncaptioned image] Transfer encoder only.
-Warm-Up and Training: [Uncaptioned image] Only decoder weights trained, [Uncaptioned image] Encoder and decoder weights trained.
-$\ddagger$: nnU-Net default (Dynamic)
+Fine-tuning strategy
+Given a pre-trained model, a crucial question arises: Which weights to transfer and how to schedule the fine-tuning?
+We investigated various different schedules, with all of themimplemented in the nnU-Net framework [18].
+Regarding weight transfer we investigated transferring (i) both the encoder and decoder [Uncaptioned image], or (ii) only the encoder [Uncaptioned image] with a randomly initialized decoder.
+Regarding the fine-tuning schedule, we investigate whether to use a learning rate warm-up of 12.5k steps, ramping the learning rate up to the maximum LR.
+When only transferring the encoder, an additional warm-up of only the decoder is investigated to adapt the randomly initialized decoder to the pre-trained encoder.
+In some configurations, this results in two learning rate warm-ups of 12.5k steps each.
+Additionally, we investigate whether to keep the encoder frozen [Uncaptioned image] for the entire fine-tuning process, or to fine-tuning encoder weights as well.
+Lastly, we investigate whether to decrease the peak learning rate to $31e-$ 3, $41e-$ 4 or to keep it at the default of $21e-$ 2.
+Results are presented in Table 3 and allow three important observations to be made: (i) Warm-up stages are essential:Not applying a warm-up step significantly reduces performance.
+Including a warm-up for both the encoder and decoder boosts accuracy by 0.6 to 1 DSC points.
+(ii) Learning rate adjustments matter:Reducing the peak learning rate to 1e-3 during fine-tuning consistently yields better results than the default 1e-2, with the best performance seen when fine-tuning both the encoder and decoder with lower learning rates.
+(iii) Freezing encoder weights is detrimental:The encoder should not remain fixed during fine-tuning.
+Allowing the encoder to be fine-tuned improves the model’s performance compared to when only the decoder is fine-tuned.
 5 Results and Discussion
-We compare our final models S3D-B and S3D-L against VoCo (Wu et al., 2024), VolumeFusion (VF) by Wang et al.
-(2023), as well as MG Models Genesis (Zhou et al., 2021).
-The baselines are pre-trained using the same framework on the same data with the same backbone and the same hyperparameters - where possible - and are scaled to fully utilize an A100 40GB GPU, optimized for 250k steps.
-We provide explicit baseline method and configuration details in Appendix A.
-Moreover, we compare against two from-scratch baselines.
-The first, ’No (Dyn.)’, represents a non-pretrained (i.e., from scratch) default nnU-Net 3D fullres architecture that was planned and preprocessed on each downstream dataset individually, potentially resulting in different architectures, data preprocessing, and spacings.
-The second from-scratch baseline, ’No (Fixed)’, is an nnU-Net training with the same plans and preprocessing as defined by our pre-training.
-The dataset-wise mean DSC and mean NSD values across our test dataset suite are provided in Table 3.
-Additionally, ranking stability of the methods is evaluated through bootstrapping and is provided in Fig.
-3.
+We compare our final model S3D against Volume Contrastive (VoCo) [42], VolumeFusion (VF) by Wang et al.
+[39], as well as Models Genesis (MG) [44].
+The baselines are pre-trained for 250k steps using the same framework on the same data with the same backbone and the same hyperparameters (where possible) but scaled to fully utilize an A100 40GB GPU.
+We provide explicit baseline methods and configuration details in Appendix A.
+Moreover, we compare against two strong from-scratch baselines.
+The first, No (Dyn.), represents a non-pretrained (i.e., from scratch) default nnU-Net 3D fullres architecture that was planned and preprocessed on each downstream dataset individually, potentially resulting in different architectures, data preprocessing, and spacings.
+The second from-scratch baseline, No (Fixed), is a nnU-Net training with the same plans and pre-processing as defined by our pre-training.
+The dataset-wise mean DSC and mean NSD values across our test dataset suite are provided in Tab.
+4.
+Additionally, the ranking stability of the methods is evaluated through bootstrapping and is provided in Fig.
+2.
 5.1 Observations
@@
 Across all tested datasets, SSL pre-trained methods demonstrate improved downstream segmentation performance.
-Comparing our S3D-B method to the most similar from-scratch baseline ’No (Fixed)’, we observe higher DSC scores in 10 out of 11 test datasets, with an average increase of +2 DSC points and +1.6 NSD points.
-This improvement is not limited to our method; MG and VF also achieve higher performance than the baseline, indicating the utility of SSL methods when applied to sufficient data and a state-of-the-art architecture.
+Comparing our S3D method to the most similar from-scratch baseline No (Fixed), we observe higher DSC scores in 10 out of 11 test datasets, with an average increase of +2 DSC points and +1.6 NSD points.
+This improvement is not limited to our method; MG and VF also achieve higher performance than the baseline, indicating the utility of SSL methods when applied to sufficient data and state-of-the-art architecture.
 MAEs dominate
-Throughout our test dataset pool, SSL schemes using the masked image modeling paradigm (MG, S3D-B, and S3D-L) consistently rank higher than the contrastive VoCo or the pseudo-segmentation-based VolumeFusion pre-training method for CNN pre-training.
-Given the age of ’Models Genesis’ - published in 2019 - it is surprising to see it outperform the more recent VoCo or VF.
+Throughout our test dataset pool, SSL schemes using the masked image modeling paradigm (MG and S3D) consistently rank higher than the contrastive VoCo or the pseudo-segmentation-based VolumeFusion pre-training method for CNN pre-training.
+Given the age of Models Genesis - published in 2019 - it is surprising to see it outperform the more recent VoCo or VF.
 We attribute this to a combination of two factors: 1.
-Models Genesis was originally published and trained on an outdated 3D-UNet (Çiçek et al., 2016) and outside of the powerful nnU-Net framework (Isensee et al., 2021).
+Models Genesis was originally published and trained on an outdated 3D-UNet [6] and outside of the powerful nnU-Net framework [18].
 This highlights the importance of avoiding Pitfall 2: Training on a state-of-the-art backbone.
@@
 VoCo and VF were introduced in conjunction with architectures they were optimized for.
-By transferring them to a CNN setting, hyperparameters chosen to optimize the method for their original architecture-pretraining combination may be suboptimal for the new CNN backbone.
-S3D-B ranks first with respect to DSC and Normalized Surface Distance (NSD), while S3D-L and MG share the second place.
-Although S3D-L and MG are very close in Average DSC, S3D-L consistently achieves lower ranks across all datasets.
-Moreover, according to the bootstrapped aggregated rank (Fig.
-3), this superior ranking is reliably higher than that of MG.
-These results indicate the overall efficacy of our pre-training method compared to currently established methods for CNN pre-training.
-SSL Method No (Dyn.) No (Fixed) VoCo VF MG S3D-B S3D-L
+By transferring them to a CNN setting, hyperparameters chosen to optimize the method for their original architecture pre-training combination may be suboptimal for the new CNN backbone.
+Table 4: S3D out-performs all baselines: The upper table presents the DSC results, while the lower table displays the NSD results across the test datasets, covering a diverse range of brain MRI tasks.
+No Fixed represents a from-scratch baseline sharing the same architecture, preprocessing and downstream training steps as all SSL methods.
+No Dyn.
+represents the original nnU-Net adapted to each downstream dataset individually.
+SSL Method No (Dyn.) No (Fix.) VoCo VF MG S3D
 Dataset Dice Similarity Coefficient (DSC)
-MS FLAIR (D1) 57.81 59.82 59.70 59.29 58.64 60.35 59.85
-Brain Mets (D2) 63.66 56.53 56.25 61.01 65.39 65.24 64.81
-Hippocampus (D3) 89.18 89.24 88.78 89.03 89.38 89.60 89.34
-Atlas22 (D4) 63.28 65.52 62.97 65.76 65.93 66.95 64.58
-CrossModa (D5) 85.64 83.44 83.07 84.24 83.91 84.08 84.02
-Cosmos22 (D6) 60.28 78.17 77.40 80.09 79.67 80.00 80.01
-ISLES22 (D7) 77.94 79.44 78.14 78.96 78.85 79.70 79.89
-Hanseg (D8) 59.00 61.85 57.47 61.49 62.52 62.11 61.93
-HNTS-MRG24 (D9) 66.73 65.90 67.65 63.34 68.00 68.62 67.94
-BRATS24 Africa (D10) 93.07 92.51 91.97 92.16 92.36 92.19 92.90
-T2 Aneurysms (D11) 46.76 41.97 40.16 44.96 45.48 47.26 44.15
-Avg. DSC 69.40 70.40 69.41 70.94 71.83 72.37 71.77
-Avg. Rank 4.64 4.55 6.27 4.36 3.18 2.00 3.00
+MS FLAIR (D1) 57.81 59.82 59.70 59.29 58.64 60.35
+Brain Mets (D2) 63.66 56.53 56.25 61.01 65.39 65.24
+Hippocampus (D3) 89.18 89.24 88.78 89.03 89.38 89.60
+Atlas22 (D4) 63.28 65.52 62.97 65.76 65.93 66.95
+CrossModa (D5) 85.64 83.44 83.07 84.24 83.91 84.08
+Cosmos22 (D6) 60.28 78.17 77.40 80.09 79.67 80.00
+ISLES22 (D7) 77.94 79.44 78.14 78.96 78.85 79.70
+Hanseg (D8) 59.00 61.85 57.47 61.49 62.52 62.11
+HNTS-MRG24 (D9) 66.73 65.90 67.65 63.34 68.00 68.62
+BRATS24 Africa (D10) 93.07 92.51 91.97 92.16 92.36 92.19
+T2 Aneurysms (D11) 46.76 41.97 40.16 44.96 45.48 47.26
+Avg. DSC 69.40 70.40 69.41 70.94 71.83 72.37
+Avg. Rank 4.64 4.55 6.27 4.36 3.18 2.00
 Dataset Normalized Surface Distance (NSD)
-MS FLAIR (D1) 78.77 80.16 79.70 79.57 79.16 80.03 80.40
-Brain Mets (D2) 80.72 76.72 72.77 79.20 81.51 82.53 82.32
-Hippocampus (D3) 99.46 99.42 99.43 99.46 99.39 99.46 99.44
-Atlas22 (D4) 70.52 73.77 70.15 73.67 74.22 75.35 73.45
-CrossModa (D5) 99.85 99.76 99.72 99.78 99.74 99.81 99.80
-Cosmos22 (D6) 72.60 96.47 94.48 96.89 96.95 97.45 96.75
-ISLES22 (D7) 88.55 90.45 89.39 90.28 89.59 90.59 90.72
-Hanseg (D8) 82.20 85.94 80.44 85.29 85.94 85.80 86.20
-HNTS-MRG24 (D9) 71.83 71.26 73.47 67.88 73.22 74.07 73.17
-BRATS24 Africa (D10) 95.66 95.36 94.95 94.94 95.33 95.06 95.72
-T2 Aneurysms (D11) 62.24 55.56 51.79 58.97 59.38 61.18 57.07
-Avg. NSD 82.04 84.08 82.39 84.17 84.95 85.58 85.00
-Avg. NSD Rank 4.27 4.27 5.82 4.64 4.00 2.18 2.82
-Table 3: S3D out-performs all baselines: Mean DSC and NSD results across the test datasets, representing a broad selection of brain MR tasks, are presented.
-’No Fixed’ represents a from-scratch baseline sharing the same architecture, preprocessing and downstream training steps as all SSL methods.
-’No Dyn.’ represents the original nnU-Net adapted to each downstream dataset individually.
-Figure 3: Dataset rank stability In addition to absolute mean performance, we report the ranking stability of the methods through bootstrapping for all test datasets as well as the aggregated rank across all datasets.
+MS FLAIR (D1) 78.77 80.16 79.70 79.57 79.16 80.03
+Brain Mets (D2) 80.72 76.72 72.77 79.20 81.51 82.53
+Hippocampus (D3) 99.46 99.42 99.43 99.46 99.39 99.46
+Atlas22 (D4) 70.52 73.77 70.15 73.67 74.22 75.35
+CrossModa (D5) 99.85 99.76 99.72 99.78 99.74 99.81
+Cosmos22 (D6) 72.60 96.47 94.48 96.89 96.95 97.45
+ISLES22 (D7) 88.55 90.45 89.39 90.28 89.59 90.59
+Hanseg (D8) 82.20 85.94 80.44 85.29 85.94 85.80
+HNTS-MRG24 (D9) 71.83 71.26 73.47 67.88 73.22 74.07
+BRATS24 Africa (D10) 95.66 95.36 94.95 94.94 95.33 95.06
+T2 Aneurysms (D11) 62.24 55.56 51.79 58.97 59.38 61.18
+Avg. NSD 82.04 84.08 82.39 84.17 84.95 85.58
+Avg. NSD Rank 4.27 4.27 5.82 4.64 4.00 2.18
+Figure 2: S3D ranks best across all methods.
+In addition to absolute mean performance, we report the ranking stability of the methods through bootstrapping for all test datasets as well as the aggregated rank across all datasets.
 Impact of dynamic configuration
-Comparing the ’No (Dyn.)’ and ’No (Fixed)’ configurations, both trained from scratch, reveals that selecting the appropriate configuration for each dataset can significantly influence performance.
+Comparing the No (Dyn.) and No (Fixed) configurations, both trained from scratch, reveals that selecting the appropriate configuration for each dataset can significantly influence performance.
 For instance, on datasets D2 and D11, the dynamic configuration outperforms the fixed by +7 and +5 DSC points, respectively, while for D6, the fixed configuration yields results +18 DSC points higher.
@@
 To measure the benefits of pre-training in such a low-data regime, we artificially reduced the total amount of data available for training to 10, 20, 30, or 40 labeled images.
-Results are presented in Table 4.
-It is evident that our pre-trained S3D-B model leads to better downstream performance compared to training from scratch in this setting.
+Results are presented in Tab.
+6.
+Our pre-trained S3D model leads to better downstream performance compared to training from scratch in this setting.
 With just 40 trained images, the fine-tuned model nearly matches the performance of the from-scratch model trained on the full dataset.
@@
 First, we evaluated fine-tuning our method on an unseen modality using the TOF Angiography Aneurysms dataset (D12).
-As shown in Table 7, without pretraining, the fixed configuration suffers a performance drop of 20 Dice points.
+As shown in Tab.
+9 without pre-training, the fixed configuration suffers a performance drop of 20 Dice points.
 We attribute this to the significant difference in median spacing for the downstream task ($[0.50,0.43,0.43]$ mm), which has a higher resolution than the fixed target spacing of $[1,1,1]$ mm used in the pre-training experiments.
@@
 Pre-training Time
-Previous studies have demonstrated the positive impact of extended training schedules on the quality of learned representations for downstream tasks (He et al., 2022; Feichtenhofer et al., 2022).
+Previous studies have demonstrated the positive impact of extended training schedules on the quality of learned representations for downstream tasks [14, 9].
 To explore the relevance of this factor in the 3D medical domain, we conducted a similar experiment, evaluating training durations ranging from 62.5k to 1M steps.
-Our results indicate that the benefits of longer training schedules begin to degrade after 250k steps, as evident in Table 5.
-This could explain why our scaled model did not achieve further performance improvements.
+Our results indicate that the benefits of longer training schedules begin to degrade after 250k steps, as evident in Tab.
+7.
 Fine-tuning length
 Initializing from pre-trained weights has the potential to reduce the computational resources needed for the network to adapt to new tasks.
-To assess this, we tested different pre-training durations on our development datasets.
+To assess this, we tested different fine-tuning durations on our development datasets.
 While maintaining 12.5k iterations for both the decoder and full network warm-up phases, we experimented with varying subsequent training lengths.
-As shown in Table 8, adding just 12.5k additional iterations (37.5k total) already outperforms training from scratch.
+As shown in Tab.
+10, adding just 12.5k additional iterations (37.5k total) already outperforms training from scratch.
 However, achieving optimal performance still requires completing the full fine-tuning schedule.
@@
 Additionally, we tested freezing the stem weights during the decoder warm-up phase.
-As shown in Table 6, the most stable and consistently effective approach was replicating the pre-trained stem and keeping it frozen during the decoder’s warm-up period.
+As shown in Tab.
+8, the most stable and consistently effective approach was replicating the pre-trained stem and keeping it frozen during the decoder’s warm-up period.
 6 Conclusion
@@
 By overcoming key pitfalls in previous research, such as limited dataset sizes, outdated architectures, and insufficient evaluation, we show a consistent performance improvement over previous SSL methods.
-Notably, for the first time, we achieve consistent gains over the dynamic, dataset-adaptive nnU-Net baseline, validated across a large and diverse set of development and testing datasets.
-While our findings are promising, several avenues remain open for future exploration.
-Notably, increasing training time and batch size did not lead to performance gains, but the question remains of whether scaling the pre-training dataset size or the model parameters could unlock new potential.
+Notably, for the first time, we achieve consistent improvements over the dynamic, dataset-adaptive nnU-Net baseline, validated across a large and diverse set of development and testing datasets.
+Limitations
+In our current manuscript, we justify our selection of baseline methods by excluding prior work that was specific to architectures (3DMAE), methods that had no public code (GL-MAE), and methods that do not outperform a nnU-Net baseline trained from scratch (see Appendix Tab.
+11).While our findings are promising, several avenues remain open for future exploration.
+We limited our method to Head and Neck MRI images, allowing us to focus on this region with a comparatively large pretraining dataset.
+Expanding to encompass the entire body across multiple imaging modalities would have required substantially more data to achieve comparable performance.
+Notably, in our ablations, we tested increasing training time and batch size (Appendix Tab.
+12) which did not lead to performance improvements, but the question remains of whether scaling the pre-training dataset size or the model parameters could enable further improvements.
 Furthermore, the intensity shifts employed by Model Genesis SSL task hint at intriguing possibilities for improving generalization across unseen MRI modalities, which needs to be further explored for MAEs.
@@
 While we applied basic filtering to exclude low-quality data, more sophisticated filtering techniques could significantly enhance the quality of the pre-training process.
-This work follows the spirit of prior studies like nnU-Net Isensee et al.
-(2021), or XXX, showing that a robust development strategy, informed model configuration, and rigorous validation lead to true and sustainable performance gains, contrasting the current hype for employing and modifying the latest network architecture.
-With our dynamic open-source framework, we hope to contribute to a cultural shift in the community towards validation-driven development enabling true scientific progress.
+To conclude, this work follows the spirit of prior studies like nnU-Net [18] showing that a robust development strategy, informed model configuration, and rigorous validation leads to true and sustainable performance improvements, contrasting the current hype for employing and modifying the latest network architecture.
+With this paper, we hope to contribute to a cultural shift in the community toward validation-driven development enabling true scientific progress.
 Appendix A Method configuration
@@
 For all baseline methods we utilize the same pre-training dataset with the same image preprocessing.
-Moreover we use the same amount of pre-training steps (250k) as for our S3D-B method and the same fine-tuning scheme, as highlighted in Table 8.
-Aside from this, we employ the SGD optimizer with LR 1e-2 with a PolyLR schedule, momentum 0.99 and weight decay 3e-5 across all pre-training experiments, as they showed to be highly robust and reliable in the supervised medical image segmentation setting using CNNs (Isensee et al., 2021).
+Moreover we use the same amount of pre-training steps (250k) as for our S3D-B method and the same fine-tuning scheme, as highlighted in Tab.
+10.
+Aside from this, we employ the SGD optimizer with LR 1e-2 with a PolyLR schedule, momentum 0.99 and weight decay 3e-5 across all pre-training experiments, as they showed to be highly robust and reliable in the supervised medical image segmentation setting using CNNs [18].
 Moreover, we denote that all these methods have their backbones replaced with a ResEnc U-Net to minimize confounding effects of different architectures.
 A.1 Models Genesis
-Models genesis (Zhou et al., 2021) pre-text task is centered around restoring original patches from transformed versions.
+Models genesis [44] pre-text task is centered around restoring original patches from transformed versions.
 The transformed version is achieved by applying four different transformations in various combinations, with the following transformations: a composition of four separate pre-training schemes: (i) Non-linear intensity transformation:Alters the intensity distribution while preserving the anatomy, focusing on learning the appearance of organs.
@@
 The entire set of hyperparameters of Models Genesis are contained within the data-augmentation.
-This allows us to transfer this transformation pipeline, as provided in the official https://github.com/MrGiovanni/ModelsGenesis without any changes to the hyperparameters.
+This allows us to transfer this transformation pipeline, as provided in the official repository without any changes to the hyperparameters.
 A.2 VolumeFusion
-Volume Fusion (Wang et al., 2023) is a pseudo-segmentation task using two sub-volumes from different 3D scans, which are fused together based on random voxel-level fusion coefficients.
+Volume Fusion [39] is a pseudo-segmentation task using two sub-volumes from different 3D scans, which are fused together based on random voxel-level fusion coefficients.
 The fused image is treated as input, and the model predicts the fusion category of each voxel, mimicking a segmentation task.
@@
 Moreover the amount of rectangles sampled is an important parameter.
-Like in the original paper we sample M Ũ(10, 40) different rectangles, iteratively.
+Like in the original paper we sample $M\sim\mathcal{U}(10,40)$ different rectangles, iteratively.
 Lastly, the number of categories was chosen to be 5, as in the original paper (this represent K = 4).
 A.3 VoCo
-The ’Volume Contrastive Learning Framework’ (VoCo) (Wu et al., 2024) is designed to enhance self-supervised learning for 3D medical image analysis by leveraging the consistent contextual positions of anatomical structures.
+The ’Volume Contrastive Learning Framework’ (VoCo) [42] is designed to enhance self-supervised learning for 3D medical image analysis by leveraging the consistent contextual positions of anatomical structures.
 The method involves generating base crops from different regions of 3D images and using these as class assignments.
@@
 Moreover we increased the target crop size from 4 originally to 5 and increased the batch size from 6 (default in our other experiments) to 12, to fully utilize the 40GB VRAM of an A100 node.
-Appendix B Additional results
+Appendix B Longer Training schedule
+MAEs are known to benefit from increasing the length of the training schedule, as shown in He et al.
+[14].
+We evaluate if this effect transfers to 3D medical pre-training by increasing the training batch size by x8 to $48$, the learning rate to $23e-$ 2, and the iteration steps by x5 to a total of $1.25M$ steps.
+We refer to this model as S3D-Long to denote the longer training schedule with more data seen.
+We denote that the architecture remains identical to the previous architecture, to isolate the effect of the length of the steps as well as the amount of samples seen.
+Results are presented in Tab.
+12.
+It can be observed that this x32 actually leads to a decrease in overall model performance on our test datasets, showing a 0.6% lower average DSC as well as a 0.6% lower Average NSD.
+The observed performance degradation of the S3D-Long model, despite the longer training schedule and increased data exposure, suggests several possible factors at play.
+While MAEs have shown benefits from extended training schedules in general computer vision tasks, the same assumptions may not directly transfer to 3D medical image pre-training due to the unique nature of this domain.
+The findings highlight the importance of tailoring training strategies to the domain.
+Table 5: Public available checkpoint trained on the ABCD dataset: To provide a public available checkpoint, we retrained our proposed model on the ABCD dataset, indicated by a *.
+It performs slightly worse than the network pre-trained on the private dataset.
+SSL Method No (Dyn.) No (Fix.) S3D* S3D
+Dataset Dice Similarity Coefficient (DSC)
+MS FLAIR (D1) 57.81 59.82 59.75 60.35
+Brain Mets (D2) 63.66 56.53 64.20 65.24
+Hippocampus (D3) 89.18 89.24 89.45 89.60
+Atlas22 (D4) 63.28 65.52 66.61 66.95
+CrossModa (D5) 85.64 83.44 83.61 84.08
+Cosmos22 (D6) 60.28 78.17 80.01 80.00
+ISLES22 (D7) 77.94 79.44 78.94 79.70
+Hanseg (D8) 59.00 61.85 61.27 62.11
+HNTS-MRG24 (D9) 66.73 65.90 67.03 68.62
+BRATS24 Africa (D10) 93.07 92.51 92.49 92.19
+Avg. DSC 71.66 73.24 74.34 74.88
+Avg. Rank 3.2 2.9 2.4 1.5
+Appendix C Additional results
 Aside from the quantitative data on the development and test dataset, we provide the quantiative data of the ablation experiments here.
 The following additional results are provided: 1.
-Results when fine-tuning in a low-data regime are presented in Table 4.
+Results when fine-tuning in a low-data regime are presented in Tab.
+6.
 2.
-Experiment on how to best transfer weights when transferring to a dataset with more than 1 input channel is provided in Table 6 3.
-Results on how the pre-training effects generalization is provided in Table 7.
+Experiment on how to best transfer weights when transferring to a dataset with more than 1 input channel is provided in Tab.
+8 3.
+Results on how the pre-training effects generalization is provided in Tab.
+9.
 4.
-Experiment results of investigating if one can reduce the fine-tuning steps are presented in Table 8
+Experiment results of investigating if one can reduce the fine-tuning steps are presented in Tab.
+10
+C.1 Public weights trained on the ABCD dataset
+Due to patient privacy concerns and data ownership regulations, we are unable to share the original pre-trained weights.
+As an alternative, we retrained our best-performing model on the National Institute of Health’s Adolescent Brain Cognitive Development (ABCD) dataset.
+This dataset comprises about 41k MRI scans with a 50-to-50 ratio of T1-weighted to T2-weighted scans.
+The results, shown in Tab.
+5, indicate that the original model slightly outperforms the version pre-trained on the ABCD dataset.
+This discrepancy is likely attributable to the greater diversity and variation in the images within our private dataset, which enables a more robust feature representation.
+C.2 Comparison to previous work on CT data
+Although this study focuses on brain MRI images, we also evaluate our method on a CT downstream task.
+Table 11 presents the results of our approach on the BTCV multi-organ segmentation task [20].
+For comparison, we incorporate a diverse set of results reported in [30], which were trained using an unspecified 80/20 data split.
+Additionally, we fine-tune the publicly available HySpark checkpoint [30] using the same five-fold cross-validation split as for our method.
+Remarkably, M3D outperforms all other methods, despite being pretrained exclusively on brain MRI data.
+Notably, none of the related approaches surpass our backbone trained on scratch.
+This highlights the critical role of leveraging state-of-the-art networks and advanced training frameworks, such as nnU-Net.
+[18, 19].
+Table 6: Forty images with SSL are almost as good as all data from-scratch!
+S3D model almost reaches the performance of the model trained from-scratch with only 40 training cases.
+* D1 has only 38 training cases for the train split.
 SSL Method N Train D1 D2 D3 D4 D5 Avg. D1-D5
@@
 full 49.37 69.13 88.78 60.74 81.33 69.87
-S3D-B (ours) 10 43.48 48.44 84.12 41.51 77.70 59.05
+S3D (ours) 10 43.48 48.44 84.12 41.51 77.70 59.05
 20 46.58 65.30 86.61 45.50 79.52 64.70
@@
 full 51.49 74.01 88.83 62.39 81.54 71.65
-Table 4: Forty images with SSL are almost as good as all data from-scratch!
-S3D-B model almost reaches the performance of the model trained from-scratch with only 40 training cases.
-* D1 has only 38 training cases for the train split.
 Table 7: Pre-training length ablation: Longer pre-training does not lead to improved performance. Interestingly, when exceeding 250k steps.
@@
 1M 50.45 71.55 88.92 62.78 81.82 71.10 448
+Table 8: Replicating the pre-trained stem weights and freezing them during the decoder warm-up phase yields the most stable and equally best results.
 Initialization Decoder Warm-Up Fold 0 Fold 1 Fold 2 Fold 3 Fold 4 Average STD
@@
 Random Unfrozen 72.20 63.16 62.25 66.71 61.47 65.16 4.42
-Table 6: Replicating the pre-trained stem weights and freezing them during the decoder warm-up phase yields the most stable and equally best results.
+Table 9: Pre-training can improve generalization: We investigate generalization to a new modality time-of-flight (ToF) MRI (top), and the generalization of a resulting method when translating it to a different clinic (bottom).
 Experiment Setting No Dyn. No Fixed VoCo VF MG S3D-B S3D-L
@@
 Patient shift Brain Mets (D13) 64.08 61.61 56.78 63.95 64.22 64.54 64.95
-Table 7: Pre-training can improve generalization: We investigate generalization to a new modality time-of-flight (ToF) MRI (top), and the generalization of a resulting method when translating it to a different clinic (bottom).
 Table 10: Fine-tuning length: When initializing from our pre-trained checkpoint, it is possible to achieve a large fraction of the final performance after less than 15% of the normal training time.
@@
 275k 53.10 71.24 89.14 63.55 82.53 71.91
+Table 11: S3D outperforms all related work on the BTCV dataset [20].
+Despite being pretrained exclusively on brain MRI data, our network outperforms all related methods.
+Values above the line are sourced from Tang et al.
+[30], which used an unknown 80/20 split.
+To fine-tuned their published pre-trained weights using a 5-fold cross-validation.
+Below the line, all models, including ours, were trained on the same 5-fold cross-validation.
+Notably, even though all related work leveraged CT data for pretraining, none surpassed the performance of our backbone model trained from scratch, emphasizing the importance of Pitfall 2.
+Pre-training Method Network Spl Kid Gall Eso Liv Sto Aor IVC Veins Pan AG Avg
+vox2vec [10] 3D UNet(FPN) [27] 91.40 90.70 59.50 72.70 96.30 83.20 91.30 83.90 69.20 73.90 65.20 79.50
+SUP [12] Swin UNETR [12] 84.20 86.70 58.40 70.40 94.40 76.00 87.70 82.10 67.00 69.80 61.00 75.80
+MAE [14] UNETR [13] 90.71 87.63 62.50 72.60 96.09 94.73 86.11 90.36 71.00 75.47 63.77 79.07
+SimMIM [33] Swin UNETR [12] 88.33 86.82 62.43 74.36 92.35 90.70 83.03 87.43 68.04 68.43 58.65 76.44
+SparK [33] MedNeXt [28] 90.92 87.66 62.43 74.36 95.03 84.85 86.04 80.63 68.83 76.57 61.43 79.21
+HySparK [30] MedNeXt+ViT  [30] 90.67 88.32 68.18 74.20 95.03 87.46 90.17 84.50 70.04 78.36 66.75 80.67
+No MedNeXt+ViT  [30] 90.35 87.46 63.18 74.49 95.09 86.00 89.29 83.22 71.85 79.48 62.59 80.27
+HySparK [30] MedNeXt+ViT [30] 90.94 86.99 63.43 74.39 95.12 87.15 88.92 83.48 72.77 79.66 64.84 80.67
+No (Dyn.) nnU-Net [18] 90.44 88.52 68.86 78.14 95.53 88.06 91.59 86.47 76.27 81.78 71.06 83.34
+No (Fix.) ResEncL (fixed) [19] 91.97 89.58 68.76 79.18 95.96 91.97 92.80 87.16 77.29 84.01 72.21 84.63
+S3D ResEncL (fixed) [19] 92.00 90.40 70.77 78.71 96.01 92.51 92.83 87.04 77.28 84.79 72.58 84.99
+Table 12: Longer training schedule degrades performance.
+When training with a larger batch size, higher learning rate, and more train steps we observe a degradation in performance for DSC and NSD.
+Ranks are calculated only between the four methods presented in the table.
+Dice Similarity Coefficient
+Dataset No Dyn. No Fixed S3D S3D-Long
+MS FLAIR (D1) 57.81 59.82 60.35 59.85
+Brain Mets (D2) 63.66 56.53 65.24 64.81
+Hippocampus (D3) 89.18 89.24 89.60 89.34
+Atlas22 (D4) 63.28 65.52 66.95 64.58
+CrossModa (D5) 85.64 83.44 84.08 84.02
+Cosmos22 (D6) 60.28 78.17 80.00 80.01
+ISLES22 (D7) 77.94 79.44 79.70 79.89
+Hanseg (D8) 59.00 61.85 62.11 61.93
+HNTS-MRG24 (D9) 66.73 65.90 68.62 67.94
+BRATS24 Africa (D10) 93.07 92.51 92.19 92.90
+T2 Aneurysms (D11) 46.76 41.97 47.26 44.15
+Avg. DSC 69.40 70.40 72.37 71.77
+Avg. Rank 3.09 3.27 1.55 2.09
+Normalized Surface Distance
+MS FLAIR (D1) 78.77 80.16 80.03 80.40
+Brain Mets (D2) 80.72 76.72 82.53 82.32
+Hippocampus (D3) 99.46 99.42 99.46 99.44
+Atlas22 (D4) 70.52 73.77 75.35 73.45
+CrossModa (D5) 99.85 99.76 99.81 99.80
+Cosmos22 (D6) 72.60 96.47 97.45 96.75
+ISLES22 (D7) 88.55 90.45 90.59 90.72
+Hanseg (D8) 82.20 85.94 85.80 86.20
+HNTS-MRG24 (D9) 71.83 71.26 74.07 73.17
+BRATS24 Africa (D10) 95.66 95.36 95.06 95.72
+T2 Aneurysms (D11) 62.24 55.56 61.18 57.07
+Avg. NSD 82.04 84.08 85.58 85.00
+Avg. NSD Rank 2.82 3.18 2.00 2.00
+Figure 3: Distribution of our pre-training dataset.
+The dataset stems from 44 centers and includes 8400 Patients with a 60 to 40 female-to-male ratio.
+Most patients were imaged with a 1.5 Tesla Philips Achieva or Ingenia scanner.
+The most prevalent modalities are T1 and T2-weighted images with some additional FLAIR images present.
+While other modalities were in the dataset, these were not used as prevalence was deemed too low.

### v3

@@
 We believe this lack of widespread adoption of previously established methods can be attributed to three major pitfalls of previous SSL research in this field:
-P1 - Limited dataset size: Many SSL approaches have been developed on very few unlabeled volumes, often being trained on fewer than 10,000 images [44, 42, 30, 31, 45, 12], almost approaching scales of supervised dataset sizes.
+P1 - Limited dataset size: Many SSL approaches have been developed on very few unlabeled volumes, often being trained on fewer than 10,000 images [48, 46, 33, 34, 49, 14, 13, 3], almost approaching scales of supervised dataset sizes.
 These datasets tend to be pooled from publicly available annotated datasets, as larger datasets pose a greater hurdle to acquire.
 While many hospitals possess 3D medical images in the millions, they are locked away from the public due to patient privacy concerns.
-Some larger open-source datasets exist, e.g., the Adolescent Brain Cognitive Decline (ABCD) dataset of the NIH (N=40k) or the UK Biobank (UK-BB) (N=120k), but they restrict access pending an internal review board’s approval, posing a hurdle for open science.
-P2 - Outdated Backbones: Many studies develop their SSL method on non-state-of-the-art architectures, e.g., utilizing transformers [42, 31, 39, 5].
+Some larger open-source datasets exist, e.g., the Adolescent Brain Cognitive Development (ABCD) dataset of the NIH (N=40k) [27] or the UK Biobank dataset (N=120k), but they restrict access pending an internal review board’s approval, posing a hurdle for open science.
+P2 - Outdated Backbones: Many studies develop their SSL method on non-state-of-the-art architectures, e.g., utilizing transformers [46, 34, 43, 6, 13, 3].
 While transformers are prevalent in the 2D natural imaging domain [9], recent architectures leveraging attention [40] have so far not been able to reach state-of-the-art performance in 3D medical segmentation.
@@
 P3 - Insufficient Evaluation: Existing methods often lack rigorous evaluation, masking the methods’ efficacy.
-This is represented through: i) Evaluating on too few datasets to show generalization of the pre-training ii) Stacking multiple contributions, e.g., novel architecture with new pre-training, which does not allow one to draw conclusions if the pre-training is effective without the architecture [39] iii) Comparing to bad from-scratch baselines, like a badly configured, outdated [6] model instead of a well-configured 3D nnU-Net CNN baseline [18].
+This is represented through: i) Evaluating too few datasets to show generalization of the pre-training [39, 3] ii) Stacking multiple contributions, e.g., novel architecture with new pre-training, which does not allow one to draw conclusions if the pre-training is effective without the architecture [43] iii) Comparing to bad from-scratch baselines, like a badly configured, outdated Çiçek et al.
+[7] UNet instead of a well-configured 3D nnU-Net CNN baseline [20].
 iv) Evaluating their method on seen data they pre-trained on.
@@
 This observation is similar to the recent study by Isensee et al.
-[19], that shows that this is also prevalent in the field of medical image segmentation when training from-scratch.
+[21], which shows that this is also prevalent in the field of medical image segmentation when training from scratch.
 Table 1: Current SSL methods are limited by three pitfalls: They are trained on less than 10K 3D images (P1), they do not use a state-of-the-art backbone for 3D image segmentation as recommended by Isensee et al.
@@
   1. 1.
-    State-of-the-Art Performance: Our final model achieves superior performance compared to existing SSL-pre-trained models using the same architecture and dataset, see Fig.
+    State-of-the-Art Performance: Our final model achieves superior performance compared to existing SSL-pre-trained models using the same architecture and pre-training dataset, see Fig.
 1.
-Notably, this is the first work to demonstrate that SSL pretraining with a fixed architecture can consistently outperform a state-of-the-art, dynamically optimized nnU-Net baseline by approximately 3 Dice points across 11 diverse downstream datasets [18].
-An associated model will be made available.
+Notably, this is the first work to demonstrate that SSL pre-training with a fixed architecture can consistently outperform a state-of-the-art, dynamically optimized nnU-Net baseline by approximately 3 Dice points across 11 diverse downstream datasets [20].
   2. 2.
@@
 1), we set up a robust development environment to design our method.
-() P1: We utilize a collection of 39k 3D MRI volumes for self-supervised pre-training, substantially larger than most datasets used in prior SSL work [44, 42, 30, 31, 45, 12].
+() P1: We utilize a collection of 39k 3D MRI volumes for self-supervised pre-training, substantially larger than most datasets used in prior SSL work [48, 46, 33, 34, 49, 14, 39].
 () P2: By using the state-of-the-art Residual Encoder U-Net CNN architecture [21], we bypass the limitations of outdated or transformer-based architectures, leveraging CNNs’ superior performance in 3D medical segmentation.
 () P3: We rigorously evaluate our model across five development and eight testing datasets, covering a wide range of downstream tasks, diverse pathologies, novel image modalities, and multi-center datasets.
-Moreover, we disentangle pre-training methodology from other contributions like novel backbones and include the strong 3D nnU-Net baseline, ensuring reliable results.
+Moreover, we disentangle the pre-training methodology from other contributions like novel backbones and include the strong 3D nnU-Net baseline, ensuring reliable results.
   3. 3.
@@
 2 Development Framework
-The goal of this paper is to develop a robust SSL pretraining method.
+The goal of this paper is to develop a robust SSL pre-training method.
 Due to the limited prior work in the 3D medical domain, many design choices need to be made.
@@
 Test Datasets
-Additionally, eight hold-out test sets were used to evaluate the efficacy of our learned representations when fine-tuning them to segment other target structures.
+Additionally, eight hold-out test datasets were used to evaluate the efficacy of our learned representations when fine-tuning them to segment other target structures.
   1. 1.
@@
 For D13 all data is used for testing, as no training is conducted.
+D12 and D13 are only used to assess generalization performance, with D13 only appearing in Appendix Tab.
+9.
 3 Related Work
@@
 Methods of the same configuration are denoted through common colors.
-(a) Sparsification: The introduction of all sparsification components showed best results, with most improvements from adding the additional densification conv layer between encoder and decoder.
+(a) Sparsification: The introduction of all sparsification components showed the best results, with most improvements coming from adding the additional densification conv layer between encoder and decoder.
 Configuration D1 D2 D3 D4 D5 Avg. D1-D5
@@
 + w/ Dens. Conv. 51.02 74.07 88.91 62.81 81.50 71.66
-(b) Masking ratio: Masking ratios between 60% and 75% worked best when choosing static ratios. Overall, a dynamic range, including higher masking ratios, performs the best.
+(b) Masking ratio: Masking ratios between 60% and 75% worked best when choosing static ratios. However, a dynamic range, including higher masking ratios, performs similarly and was chosen.
 Mask ratio D1 D2 D3 D4 D5 Avg. D1-D5
@@
 (c) Development performance: When comparing our final S3D model against the baseline methods, trained equally, S3D exceeds all baselines on the development datasets.
-SSL Method No (Dyn.) No (Fix.) S3D* S3D
-Dataset Dice Similarity Coefficient (DSC)
-MS FLAIR (D1) 57.81 59.82 59.75 60.35
-Brain Mets (D2) 63.66 56.53 64.20 65.24
-Hippocampus (D3) 89.18 89.24 89.45 89.60
-Atlas22 (D4) 63.28 65.52 66.61 66.95
-CrossModa (D5) 85.64 83.44 83.61 84.08
-Cosmos22 (D6) 60.28 78.17 80.01 80.00
-ISLES22 (D7) 77.94 79.44 78.94 79.70
-Hanseg (D8) 59.00 61.85 61.27 62.11
-HNTS-MRG24 (D9) 66.73 65.90 67.03 68.62
-BRATS24 Africa (D10) 93.07 92.51 92.49 92.19
-Avg. DSC 71.66 73.24 74.34 74.88
-Avg. Rank 3.2 2.9 2.4 1.5
+Pre-training D1 D2 D3 D4 D5 Avg. D1-D5
+No Dyn. 45.56 72.26 88.80 60.44 82.61 69.93
+No Fixed 49.37 69.13 88.78 60.74 81.33 69.87
+VoCo 50.35 67.20 88.22 57.82 80.29 68.77
+VF 49.93 69.58 88.83 61.75 81.48 70.31
+MG 50.50 71.14 88.83 63.29 82.15 71.18
+S3D (ours) 51.49 74.01 88.83 62.39 81.54 71.65
 Masked autoencoders (MAEs) are a well-established pre-training paradigm in the natural imaging domain and in the medical image segmentation domain for transformers.
@@
 As a sampling strategy, we follow random masking, as previous work showed no benefit of structured masking for images nor videos [16, 10].
-In the scope of the development phase we explore 5 static masking ratios between 30 and 90%, and evaluate a dynamic masking ratio randomly masking between 60% and 90%.
-Results are presented in Table Tab.
+In the scope of the development phase, we explore 5 static masking ratios between 30% and 90% and evaluate a dynamic masking ratio randomly masking between 60% and 90%.
+Results are presented in Tab.
 2(b) and highlight that the masking ratios of 60%, 75%, and the dynamic masking ratio of 60% to 90% perform equally well.
-Due to the highly similar performance, we choose to proceed with a dynamic masking ratio over the static masking ratio, due to expecting this masking to be more difficult to learn and with the upcoming scaling experiment in mind.
+Due to the highly similar performance, we choose to proceed with a dynamic masking ratio over the static masking ratio, due to expecting this masking to be more difficult to learn.
 We refer to this model as Spark3D (S3D) going forward.
@@
 In some configurations, this results in two learning rate warm-ups of 12.5k steps each.
-Additionally, we investigate whether to keep the encoder frozen [Uncaptioned image] for the entire fine-tuning process, or to fine-tuning encoder weights as well.
-Lastly, we investigate whether to decrease the peak learning rate to $31e-$ 3, $41e-$ 4 or to keep it at the default of $21e-$ 2.
+Additionally, we investigate whether to keep the encoder frozen [Uncaptioned image] for the warm-up process or to fine-tune the encoder weights as well.
+Lastly, we investigate whether to decrease the peak learning rate to 1e-3, 1e-4 or to keep it at the default of 1e-2.
 Results are presented in Table 3 and allow three important observations to be made: (i) Warm-up stages are essential:Not applying a warm-up step significantly reduces performance.
@@
 (iii) Freezing encoder weights is detrimental:The encoder should not remain fixed during fine-tuning.
-Allowing the encoder to be fine-tuned improves the model’s performance compared to when only the decoder is fine-tuned.
+Allowing the encoder to be fine-tuned improves the model’s performance compared to when only the decoder is fine-tuned, as observed in early exploratory work.
 5 Results and Discussion
-We compare our final model S3D against Volume Contrastive (VoCo) [42], VolumeFusion (VF) by Wang et al.
-[39], as well as Models Genesis (MG) [44].
+We compare our final model S3D against Volume Contrastive (VoCo) [46], VolumeFusion (VF) [43], as well as Models Genesis (MG) [48].
 The baselines are pre-trained for 250k steps using the same framework on the same data with the same backbone and the same hyperparameters (where possible) but scaled to fully utilize an A100 40GB GPU.
@@
 Moreover, we compare against two strong from-scratch baselines.
-The first, No (Dyn.), represents a non-pretrained (i.e., from scratch) default nnU-Net 3D fullres architecture that was planned and preprocessed on each downstream dataset individually, potentially resulting in different architectures, data preprocessing, and spacings.
+The first, No (Dyn.), represents a from scratch default nnU-Net 3D architecture that was planned and preprocessed on each downstream dataset individually, potentially resulting in different architectures, data preprocessing, and spacings.
 The second from-scratch baseline, No (Fixed), is a nnU-Net training with the same plans and pre-processing as defined by our pre-training.
-The dataset-wise mean DSC and mean NSD values across our test dataset suite are provided in Tab.
+The dataset-wise mean dice similarity coefficient (DSC) and mean normalized surface distance (NSD) at 2 mm tolerance values across our test dataset suite are provided in Tab.
 4.
-Additionally, the ranking stability of the methods is evaluated through bootstrapping and is provided in Fig.
+Additionally, the ranking stability is evaluated through bootstrapping and is provided in Fig.
 2.
@@
 Comparing our S3D method to the most similar from-scratch baseline No (Fixed), we observe higher DSC scores in 10 out of 11 test datasets, with an average increase of +2 DSC points and +1.6 NSD points.
-This improvement is not limited to our method; MG and VF also achieve higher performance than the baseline, indicating the utility of SSL methods when applied to sufficient data and state-of-the-art architecture.
-MAEs dominate
-Throughout our test dataset pool, SSL schemes using the masked image modeling paradigm (MG and S3D) consistently rank higher than the contrastive VoCo or the pseudo-segmentation-based VolumeFusion pre-training method for CNN pre-training.
+This improvement is not limited to our method; MG and VF also achieve higher performance than the baseline, indicating the utility of SSL methods when applied to sufficient data and a state-of-the-art architecture.
+MAEs dominate Throughout our test dataset pool, SSL schemes using the masked image modeling paradigm (MG and S3D) consistently rank higher than the contrastive VoCo or the pseudo-segmentation-based VolumeFusion pre-training method for CNN pre-training.
 Given the age of Models Genesis - published in 2019 - it is surprising to see it outperform the more recent VoCo or VF.
@@
 Despite this decrease, pre-training mitigates some of this degradation and proves highly beneficial compared to training the same configuration from scratch.
-Interestingly, ModelGenesis achieves the best results, potentially due to its use of intensity augmentations during pre-training, which increases robustness against brightness shifts, such as when generalizing to different MRI sequences.
+Interestingly, Models Genesis achieves the best results, potentially due to its use of intensity augmentations during pre-training, which increases robustness against brightness shifts, such as when generalizing to different MRI sequences.
 Second, we fine-tuned on the D2 dataset using only the T1 contrast-enhanced (T1ce) sequence and applied these models directly to D13 without any additional fine-tuning.
-While the dynamic configuration performed best on the in-distribution validation cases of D2, the results on D13 indicate that MAE pre-training improves generalization across different centers, with our S3D model yielding the best performance.
+While the dynamic configuration performed best on the in-distribution validation cases of D2, the results on D13 indicate that pre-training improves generalization across different centers, with our S3D model yielding the best performance.
 Pre-training Time
@@
 In our current manuscript, we justify our selection of baseline methods by excluding prior work that was specific to architectures (3DMAE), methods that had no public code (GL-MAE), and methods that do not outperform a nnU-Net baseline trained from scratch (see Appendix Tab.
-11).While our findings are promising, several avenues remain open for future exploration.
+11).
+While our findings are promising, several avenues remain open for future exploration.
 We limited our method to Head and Neck MRI images, allowing us to focus on this region with a comparatively large pretraining dataset.
@@
 With this paper, we hope to contribute to a cultural shift in the community toward validation-driven development enabling true scientific progress.
+7 Acknowledgments
+This work was partly funded by the Helmholtz Foundation Model Initiative (HFMI) under the subproject ‘The Human Radiome Project’ (THRP).
+Part of this work was funded by Helmholtz Imaging (HI), a platform of the Helmholtz Incubator on Information and Data Science.
 Appendix A Method configuration
@@
 The findings highlight the importance of tailoring training strategies to the domain.
-Table 5: Public available checkpoint trained on the ABCD dataset: To provide a public available checkpoint, we retrained our proposed model on the ABCD dataset, indicated by a *.
+Table 5: Publicly available checkpoint trained on the ABCD dataset: To provide a public available checkpoint, we retrained our proposed model on the ABCD dataset, indicated by a *.
 It performs slightly worse than the network pre-trained on the private dataset.
@@
 Table 6: Forty images with SSL are almost as good as all data from-scratch!
-S3D model almost reaches the performance of the model trained from-scratch with only 40 training cases.
+The pre-trained S3D model almost reaches the performance of the model trained from-scratch with only 40 training cases, with the exception of D4.
+Overall train/val/test dataset size was 38/10/12 for D1, 67/17/21 for D2, 166/42/52 for D3, 419/105/131 for D4, 134/34/42 for D5.
+Results in the table are reported on the validation set.
+full: Uses all train samples of the dataset.
 * D1 has only 38 training cases for the train split.
@@
 Table 9: Pre-training can improve generalization: We investigate generalization to a new modality time-of-flight (ToF) MRI (top), and the generalization of a resulting method when translating it to a different clinic (bottom).
-Experiment Setting No Dyn. No Fixed VoCo VF MG S3D-B S3D-L
-Modality shift TOF Angio. Aneurysms(D12) 42.61 22.76 22.32 31.21 34.60 28.72 26.90
-In Distribution Brain Mets (D2) 72.81 67.93 64.34 71.69 69.05 71.56 72.53
-Patient shift Brain Mets (D13) 64.08 61.61 56.78 63.95 64.22 64.54 64.95
+Experiment Setting No Dyn. No Fixed VoCo VF MG S3D-B
+Modality shift TOF Angio. Aneurysms(D12) 42.61 22.76 22.32 31.21 34.60 28.72
+In Distribution Brain Mets (D2) 72.81 67.93 64.34 71.69 69.05 71.56
+Patient shift Brain Mets (D13) 64.08 61.61 56.78 63.95 64.22 64.54
 Table 10: Fine-tuning length: When initializing from our pre-trained checkpoint, it is possible to achieve a large fraction of the final performance after less than 15% of the normal training time.
 Despite this a full training schedule reaches better performance.
+These experiments were conducted using S3D long on the validation splits.
 FT Iterations D1 D2 D3 D4 D5 Avg. D1-D5
@@
 While other modalities were in the dataset, these were not used as prevalence was deemed too low.
+Appendix D Distinction to AMAEs
+Concurrently with this work, [25] introduced the AMAEs framework.
+Like our approach, it utilizes a dataset of approximately 40k 3D images and employs a state-of-the-art CNN architecture.
+However, in contrast to our more comprehensive evaluation, their assessment is based on three in-distribution and one out-of-distribution downstream dataset.
+While evaluation on three in and one out-of-distribution datasets can suffice to draw some insights, their evaluation setup has additional limitations.
+First, they constrain themselves to a low-data regime and do not assess whether their performance surpasses a default nnU-Net baseline that is trained for 1000 epochs.
+Second, their fine-tuning strategy is fixed to a single-channel input to align with the pre-trained stem weights.
+While this would be okay for their method, they apply the same limitation to their baselines, which typically utilize all available modalities, which can lead to stronger performance.
+For instance, in their evaluation on the BraTS dataset, only one of the four available modalities was used, potentially limiting the effectiveness of the baseline models.
+These choices may impact result reliability and align with Pitfall 3—unreliable evaluation practices.
+To address such concerns, our work adopts a very comprehensive evaluation strategy to allow drawing reliable conclusions on the state-of-the-art in self-supervised learning for 3D medical image segmentation for the first time.
+Appendix E Acknowledgement of ABCD
+Data used in the preparation of this article were obtained from the Adolescent Brain Cognitive DevelopmentSM (ABCD) Study (https://abcdstudy.org), held in the NIMH Data Archive (NDA).
+This is a multisite, longitudinal study designed to recruit more than 10,000 children age 9-10 and follow them over 10 years into early adulthood.
+The ABCD Study® is supported by the National Institutes of Health and additional federal partners under award numbers U01DA041048, U01DA050989, U01DA051016, U01DA041022, U01DA051018, U01DA051037, U01DA050987, U01DA041174, U01DA041106, U01DA041117, U01DA041028, U01DA041134, U01DA050988, U01DA051039, U01DA041156, U01DA041025, U01DA041120, U01DA051038, U01DA041148, U01DA041093, U01DA041089, U24DA041123, U24DA041147.
+A full list of supporters is available at https://abcdstudy.org/federal-partners.html.
+A listing of participating sites and a complete listing of the study investigators can be found at https://abcdstudy.org/consortium_members/.
+ABCD consortium investigators designed and implemented the study and/or provided data but did not necessarily participate in the analysis or writing of this report.
+This manuscript reflects the views of the authors and may not reflect the opinions or views of the NIH or ABCD consortium investigators.